### v1

# Can Large Language Models Address Open-Target Stance Detection?

Abstract
Stance detection (SD) assesses a text’s position towards a target, typically labeled as "favor," "against," or "neutral." We introduce Open-Target Stance Detection (OTSD), where targets are neither seen during training nor provided as input. Evaluating Large Language Models (LLMs) like GPT-3.5, Llama 3, and Mistral, we compare their performance with the Target-Stance Extraction (TSE) approach, which has the advantage of using predefined targets. LLMs perform better than TSE in target generation when the real target is explicitly and not explicitly mentioned in the text. For stance detection, LLMs perform better in explicit scenarios but fail in non-explicit ones. [^1] We will release the code and the dataset.
Can Large Language Models Address Open-Target Stance Detection?
Abu Ubaida Akash  and Ahmed Fahmy  and Amine Trabelsi {akaa2803, mosa2801, amine.trabelsi}@usherbrooke.ca Université de Sherbrooke
1 Introduction
Stance detection (SD) is the task of determining the position of a text or person towards a certain target, often split into “favor” and “against” with an optional third split being “neutral”, “none”, or “neither” (Sutter et al., 2024). The target word can be mentioned in the text explicitly, or sometimes the idea of the target word can be conveyed indirectly (Küçük and Can, 2020). In Zero-shot Stance Detection (ZSSD), the model predicts the stance for targets that it has not seen during the training phase. This task is essential because collecting training data for every possible topic or target is impractical (Allaway et al., 2021). While recent research has focused on ZSSD (Zhang et al., 2023; Li et al., 2023b; Wen and Hauptmann, 2023; Liang et al., 2022a; Zhu et al., 2022; Luo et al., 2022; Xu et al., 2022; Liang et al., 2022b; Li and Yuan, 2022), most studies assume that the target is known or manually identified and given as input, a situation that is uncommon in real-world applications. In practice, the target is unknown, and it is not necessarily mentioned explicitly in the text. Additionally, target annotation is a tedious and expensive task in SD (Küçük and Can, 2020; Sutter et al., 2024).
In this paper, we focus on a different yet challenging task we refer to as Open-Target Stance Detection (OTSD). In this task, the target is neither seen during training (zero-shot) nor provided as input to the model. In OTSD, one important challenge lies in identifying the target from the text rather than having it provided as input. The target may be mentioned explicitly or not. The stance is predicted with respect to the produced target.
There has been one notable attempt to partially address this real-world setting of SD. Li et al. (2023a) generate targets from the input text and later detect stance based on these targets. However, before starting the SD task, they map these generated targets to a predefined list of golden targets (details in §2) to ensure exact wording and facilitate evaluation. They called their approach Target-Stance Extraction in the Wild (TSE). This approach does not fully align with our OTSD setting, which assumes no input target information is given during the whole process, and is impractical due to the need for a comprehensive list of all possible targets. Moreover, our experiment reveals a performance gap in Li et al. (2023a)’s approach when used on text with explicit or non-explicit target mentions. This aspect was not highlighted by the authors and warrants further investigation in the context of OTSD.
In this work, we present and explore the OTSD task, focusing on its two main steps: Target Generation and Stance Detection (see Eq. 2). As a zero-shot task for both target and stance, Large Language Models (LLMs) appear well-suited for this challenge. Therefore, we examine the performance of LLMs and compare it to TSE, the primary existing work by Li et al. (2023a), which benefits from using a predefined list of targets during the process (see Eq. 1). Our empirical study aims to address the following research questions: [RQ#1] How do proprietary and open LLMs—specifically GPT-3.5 (Brown et al., 2020), Llama 3 (Touvron et al., 2023), and Mistral (Jiang et al., 2023)—perform in open target generation compared to the TSE approach when the real target is explicitly or not explicitly mentioned in the text for stance classification, and how do they compare to each other? [RQ#2] How do the same LLMs perform compared to TSE in stance detection (in the context of OTSD) for both explicit and non-explicit cases, and how do they compare to each other?
Our contributions consist of introducing the task of OTSD and conducting experiments and answer these research questions.
2 Target-Stance Extraction (TSE)
In TSE, Li et al. (2023a) demonstrate target extraction and stance detection tasks in both in-target and zero-shot settings. For our comparison, we only consider their zero-shot setting. In TSE zero-shot, they first generate the target using keyphrase generation models, map the generated targets (multiple targets per sample) to a predefined list of golden targets to find the most similar one, and finally use it to detect the stance of the text toward the target (Eq. 1). The best match from their keyphrase model’s output, marked as TSE-BestGen, is used for our comparison. The best mapped target (TSE-mapped) is considered the predicted target. For example, given the text "Embracing different faiths teaches us empathy and respect. Let’s celebrate our diverse beliefs and foster unity. #Faith #Unity #Respect", TSE primarily generates targets such as Peace, Religious diversity, Respect, etc.. From a predefined list of possible targets (e.g., Face Mask, Atheism, Donald Trump, etc.), Atheism matches most closely with Religious diversity and is considered as the final predicted target. In OTSD, we aim at directly generating Religious diversity or a very semantically similar phrasing, avoiding the constraints of a predefined list. In the second step, they detect the stance toward their generated target.
3 Open-Target Stance Detection (OTSD)
3.1 Task Definition
Given an input text $x$, previous work (Eq. 1) generates a target $t^{\prime}$ and maps it to a predefined list of targets $k$, resulting in a mapped target $t$. Finally, they detect the stance $y$ given $x$ and $t$. OTSD task (Eq. 2) is to detect the stance $y$ from the input text $x$ and the generated target $t^{\prime}$, eliminating the need for a predefined list $k$.
[…]
3.2 Approach
To address the research questions outlined in §1, our approach to OTSD leverages the zero-shot learning capabilities of LLMs. We adopt the "Task Definition" prompting strategy for the OTSD task, as it has been shown by Cruickshank and Ng (2023) to outperform other strategies—such as "Task-only", "Context Analyze", "Context Question", "Zero-shot CoT", and "CoDA"—particularly for smaller models (7B-8B parameters) across most datasets. Given that this is an open-target task, we aim to use a straightforward prompting approach to maintain fairness and generalizability; we do not tailor the prompt for specific targets or include demonstrations or in-context examples, as this remains a zero-shot task.
Since LLMs are sensitive to the length of generated text, we limit the target length to ensure a fair comparison with TSE in terms of target generation quality. To determine the optimal length, we test various target length instructions across all LLMs (§4.1) on 100 randomly selected samples from our dataset (§4.1), aiming to achieve an average length and standard deviation that closely match those of the golden targets. Based on our empirical study, we set the maximum target length to 5 for GPT-3.5 and 4 for Llama and Mistral to obtain an average length and standard deviations close to those of gold targets (explicit: $3.78\pm 1.5$; non-explicit 2.66±1.76)2.66$\pm$ 1.76)2.66 ± 1.76 ). Finally, we employ two different prompting approaches for target generation (TG) and stance detection (SD).
TG+SD (Two-Step Approach)
In this approach, we first generate the target and then sequentially detect the stance of the text toward the generated target. By directing the LLM to focus on one task at a time, this method may reduce the cognitive load on the model.
TG&SD (Single-Step Joint Approach)
This approach combines target generation and stance detection into a single step using a unified prompt. By jointly generating the target and stance, the model can develop a more holistic understanding of the relationship between the text, target, and stance.
The prompts and their design justifications can be found in Appendix A.
4 Experiments
In the OTSD task, we compare our approaches with TSE in both TG and SD. To ensure sensitivity and robustness, we run our experiments three times independently and report the average results  [^2] Score variations are negligible, up to four decimal places.. Since OTSD is a generative task, we primarily compare our results with TSE-BestGen, despite its reliance on a predefined list of golden targets. Additionally, we consider TSE-mapped as the current best score, regardless of task type. Details about our hardware settings can be found in Appendix D.
4.1 Dataset and Model
We use the same dataset as TSE’s zero-shot setup for a fair comparison, consisting of a total of 3,000 samples considering three stance classes i.e., Favor, Against, and None. This includes 500 samples each from 6 different targets (see Appendix E). Additionally, we categorize the dataset into explicit and non-explicit samples. This categorization involves removing stop words and special characters from both texts and targets, lemmatizing all words, and subsequently comparing the text with the target. A sample is labeled as explicit if any target words are found in the corresponding text; otherwise, it is labeled as non-explicit. This process results in 1,804 explicit samples and 1,196 non-explicit samples (in Appendix E). It is important to note that because we compare our approach with TSE zero-shot, and they do not provided the code or sufficient details for reproducing the generation or selection of candidate targets, we are unable to use other stance detection datasets for benchmarking.
In our focused contribution, we utilize both proprietary and non-proprietary LLMs, including GPT-3.5 (Brown et al., 2020), Llama-3 (Touvron et al., 2023), and Mistral (Jiang et al., 2023), to address our two research questions (§1). [^3] Although we do not claim that the results or findings are true for all LLMs, experimenting with other proprietary LLMs is financially constraining for us. Model descriptions can be found in Appendix  B.
4.2 Evaluation Methods
To measure the generated target quality in OTSD task, we consider the following metrics:
BTSD: We fine-tune the BERTweet model (Nguyen et al., 2020) following the same setup as TSE (fine-tuning details in Appendix F). The SD score (F1-macro) is used as the evaluator for the generated targets. To assess the effectiveness of this metric, we experiment with different levels of target quality by modifying or removing words from the golden targets, substituting them with incorrect targets, and selecting random words from the vocabulary as targets. We find that the BERTweet F1 score consistently and significantly improved with higher-quality targets, in both explicit and non-explicit cases (Appendix C, Table 3). Therefore, the F1 score on stance detection serves as a reliable proxy for evaluating target generation quality in our dataset.
Human Evaluation: To evaluate the quality of each generated target and to further validate the BTSD metric, we conduct a small-scale human evaluation using 500 randomly selected samples (300 explicit and 200 non-explicit, with 16.67% from each stance label). Three independent domain experts are asked to assess the relevance of the generated targets by evaluating its relevance to the golden target in the tweet. Each annotator is provided with an individual file, ensuring that their evaluations remain independent of one another. The full agreement details and score distribution are included in Appendix I and J, respectively.
SemSim: In OTSD, since we do not rely on a predefined list of targets, the generated targets may not precisely match the gold targets but are semantically related. Therefore, we measure the semantic similarity between the generated and gold targets using BERT embeddings. Detailed steps can be found in Appendix G.
We use macro-average F1 for the task of SD.
TG+SD TG&SD
Model SemSim BTSD HEval SemSim BTSD HEval
Explicit
TSE-M 0.96 36.63 0.716 - - -
TSE-B 0.86 35.8 0.338 - - -
GPT-3.5 0.87 38.43 0.546 0.87 39.6 0.663
Llama-3 0.88 38.31 0.54 0.87 37.75 0.558
Mistral 0.85 38.47 0.586 0.87 39.02 0.65
Non-explicit
TSE-M 0.9 30.56 0.391 - - -
TSE-B 0.82 29.32 0.215 - - -
GPT-3.5 0.84 33.1 0.419 0.83 31.32 0.405
Llama-3 0.85 33.06 0.447 0.83 33.81 0.49
Mistral 0.83 30.18 0.386 0.83 34.12 0.495
Table 1: Comparison of TG performance between TSE and LLMs across various evaluation metrics. HEval represents human evaluation. TSE-M and TSE-B correspond to TSE-mapped and TSE-BestGen, respectively. The highest score for each metric across all models is highlighted in bold, while for BTSD and HEval, the better results between the two approaches are underlined.
5 Result Analysis
5.1 Target Generation
Following RQ#1, Table 1 shows that for the explicit case, Mistral achieves the highest BTSD score in the TG+SD setting, while GPT-3.5 leads in TG&SD. Overall, LLMs outperform TSE-BestGen in the BTSD metric and produce more relevant targets according to HEval scores. Although TSE-mapped has the highest SemSim and HEval scores, its BTSD score is lower due to generating either a perfect match or something entirely different (as illustrated in Figure 1). In contrast, the LLMs generate targets that are either highly related or partially related to the golden ones, with fewer irrelevant targets compared to TSE-mapped, leading to higher BTSD scores despite lower HEval. For non-explicit cases, LLMs outperform TSE-BestGen in both BTSD and HEval, with Llama-3 generating the most relevant targets (as indicated by the HEval score), even surpassing TSE-mapped. In the explicit case, TG&SD generally outperforms TG+SD across most LLMs, except for Llama, which performs better with TG+SD in BTSD. In non-explicit cases, GPT excels with TG+SD, while the other LLMs perform better with TG&SD. Among LLMs, GPT performs best overall in the explicit case, while Mistral leads in the non-explicit case. A manual inspection (on GPT-3.5’s 30 incorrect predictions) investigating the lower score in non-explicit cases reveals that the model struggles without sufficient context, often due to the way the TSE test set is constructed. This lack of context, such as implicit hints on the target or the context of other messages in the thread, causes the model to generate targets based solely on the provided text (example in Appendix H.1), leading to inaccuracies. Table 1 also confirms that higher HEval scores correlate with higher BTSD scores, validating BTSD as a reliable proxy for evaluating target quality.
Model TG+SD TG&SD
Explicit
TSE-mapped 38.1 -
TSE-BestGen 37.81 -
GPT-3.5 42.68 47.61
Llama-3 43.13 42.1
Mistral 46.15 46.32
Non-explicit
TSE-mapped 32 -
TSE-BestGen 31 -
GPT-3.5 32.66 33.94
Llama-3 33.36 31.9
Mistral 33.61 30.47
Table 2: Comparison of SD performance between TSE and LLMs, measured by F1-macro (in %). The highest scores across all models are highlighted in bold, with the best scores between the two approaches underlined.
5.2 Stance Detection
To address RQ#2, Table 2 shows that in stance detection, LLMs without a predefined list outperform TSE, which uses a predefined list, in both approaches (TG+SD and TG&SD) for explicit cases. In non-explicit cases, the scores are comparable in the TG+SD setting, but higher scores are observed for LLMs in the TG&SD setting. This is because, in explicit cases, LLMs generate better-quality targets than TSE-BestGen, leading to improved stance detection as evidenced in Table 1. For non-explicit cases, the holistic view provided by the TG&SD setting makes it easier for LLMs to infer stance from the text. Note that although LLM generated targets are better quality than TSE-BestGen, their stance labels can still be different from golden labels, resulting in low F1 scores. Our manual analysis revealed that these errors mainly arise from insufficient context in the input text, both explicit and non-explicit (see Appendix H.2 for examples). In both approaches, LLMs achieve much higher SD scores in explicit cases compared to non-explicit ones. Again, this is due to the incorrectly generated targets in the non-explicit case revealed by manual inspection. Incorrect targets correspond in a high percentage to incorrect stance labels. Most of targets were either completely incorrect or, in a few cases, contradictory to the real target, altering the stance.
6 Conclusion
LLMs outperform TSE in target generation for both explicit and non-explicit cases, despite TSE’s use of a predefined list. However, LLMs struggle more with generating targets when they are not explicitly mentioned in the text. LLMs also perform better in stance detection than TSE for both explicit and non-explicit cases. Additionally, LLMs show better stance detection performance when the target is explicitly mentioned compared to non-explicit cases. Finally, the single-step joint prompt approach (TG&SD) is more effective than the two-step approach (TG+SD) for both TG and SD in both explicit and non-explicit scenarios.
Limitation
We do not provide a specific coherence measure to evaluate the Open-Target Stance Detection task, as the target and the stance toward it are inherently interconnected. Due to resource constraints, our experiments were conducted on a small scale in terms of the models used. The human evaluation of target quality was also limited to 500 samples. Regarding potential data leakage during the pre-training of LLMs, it is possible that these models were exposed to some of our test data, leading to prior knowledge of certain entities. However, since the LLMs were not explicitly pre-trained on the same task, we continue to refer to our approach as zero-shot stance detection.
Appendix A Prompting Details
TG+SD (Two-Step Approach)
Prompt (TG):
You will be provided with a tweet, and your task is to generate a target for this tweet. A target should be the topic on which the tweet is talking. The target can be a single word or a phrase, but its maximum length MUST be 4 words. The output should only be the target, no other words. Do not provide any explanation but you MUST give an output, do not leave any output blank.
Prompt (SD):
Stance classification is the task of determining the expressed or implied opinion, or stance, of a statement toward a certain, specified target. Analyze the following tweet and determine its stance towards the provided target. If the stance is in favor of the target, write FAVOR, if it is against the target write AGAINST and if it is ambiguous, write NONE. Do not provide any explanation but you MUST give an output, do not leave any output blank. Only return the stance as a single word, and no other text.
Prompt Justification:
  • Clarity of Task: The task description explicitly states what the LLM needs to do—generate a target for a given tweet in TG prompt and for SD, it provides a clear and concise definition of stance classification. This ensures that the LLM understands the primary objective without ambiguity.
  • Focus on Relevance: By instructing the model that the target should be the topic of the tweet, it emphasizes the need for relevance, ensuring that the generated target accurately represents the tweet’s content.
  • Conciseness Requirement: Limiting the target to a maximum of 4 words ensures that the output remains concise and focused, which is crucial for the generated target to be compared fairly with the golden target.
  • Stance Options Provided: By specifying the possible stances (Favor, Against, None), the prompt sets clear categories for the model to choose from. This simplifies the models decision process and ensures the outputs are standardized.
  • Output Format Specification: Clearly specifying that the output should only be the target, with no additional words or explanations, helps to standardize the outputs, making them easier to evaluate.
  • Mandating Output Generation: By instructing the model not to leave any output blank and always to provide an answer, the prompt ensures robustness in responses, reducing the chance of missing data.
TG&SD (Single-Step Joint Approach)
Prompt (TG, SD):
Stance classification is the task of determining the expressed or implied opinion, or stance, of a statement toward a certain, specified target. Analyze the following tweet, generate the target for this tweet, and determine its stance towards the generated target. A target should be the topic on which the tweet is talking. The target can be a single word or a phrase, but its maximum length MUST be 4 words. If the stance is in favor of the target, write FAVOR, if it is against the target write AGAINST and if it is ambiguous, write NONE. If the stance is in favor of the generated target, write FAVOR, if it is against the target write AGAINST and if it is ambiguous, write NONE. The answer only has to be one of these three words: FAVOR, AGAINST, or NONE. Do not provide any explanation but you MUST give an output, do not leave any output blank. The output format should be: “‘Target: <target>, Stance: <stance>“‘.
Prompt Justification:
  • Integrated Task Definition: The prompt combines the definitions of both target generation and stance classification, ensuring the LLM understands the combined objective of generating a target and determining the stance towards it.
  • Clear and Concise Instructions: By clearly outlining the steps—first generating a target and then determining the stance towards it—the prompt guides the model through a logical sequence, reducing ambiguity and improving task performance.
  • Conciseness Requirement: Limiting the target to a maximum of 4 words ensures that the output remains concise and focused, which is crucial for the generated target to be compared fairly with the golden target.
  • Target Generation Constraints: Limiting the target to a maximum of 4 words ensures that the output remains concise and focused, which is crucial for the generated target to be compared fairly with the golden target.
  • Stance Classification Options: Providing explicit stance options (Favor, Against, None) simplifies the decision process for the model and ensures consistency in the outputs.
  • Mandatory Output Generation: By instructing the model to always provide an output and not leave any fields blank, the prompt ensures robustness and completeness in the responses, reducing the chance of missing data.
  • Output Format Specification: Specifying the output format as ’Target: <target>, Stance: <stance>’ standardizes the responses, making them easier to parse and evaluate.
  • Focus on Simplicity and Relevance: The prompt emphasizes that only the stance should be returned as a single word and no additional text, which maintains simplicity and relevance in the outputs.
Appendix B Model Descriptions
GPT-3.5-turbo-0125
is a variant of the GPT-3.5 series with 175 billion parameters which utilizes a transformer-based architecture. This model is decoder-only, similar to models like GPT-3 (Brown et al., 2020) and GPT-4 (Achiam et al., 2023), but with optimizations for efficiency and performance.
Llama-3-8B Instruct
is a decoder-only model in the LLaMA (Touvron et al., 2023) series with 8 billion parameters that follows the architectural principles of its predecessor, Llama-2 (Touvron et al., 2023), but has been instruction-tuned to improve its performance on a variety of tasks.
Mistral 7B Instruct
(Jiang et al., 2023) is a decoder-only model like Llama-2 (Touvron et al., 2023), but with different architectural features like Grouped-Query Attention and Sliding Window Attention. The Mistral model is 7 billion parameters and has been instruction tuned.
Appendix C BTSD Metric Justification
Input F1 (macro)
Explicit Non-explicit
Tweet only 31.02 30.37
-w/ Gold Target 41.67 34.65
-w/ altered GT 35.3 32.54
-w/ incorrect Target 25.58 23.58
-w/ random vocab 18.66 14
Table 3: Macro avg. F1-Score of BertTweet on Stance Detection with different quality levels of the input target.
Table 3 contains the results of the stratified experiments with different measure stance detection F1 score when giving as input the text and a target, varying the quality of this latter. The numbers are significantly different.
Appendix D Hardware Setting
All of the experiments were run on a computer with Ubuntu 22.04 Linux with x64 CPU with 4 cores, 32 GB of RAM and one NVIDIA V100 GPU.
Appendix E Dataset Samples
Targets: Creationism (Somasundaran and Wiebe, 2010), Gay Rights (Somasundaran and Wiebe, 2010), Climate Change is a Concern (Mohammad et al., 2016), MeToo Movement (Gautam et al., 2020), Merger of Disney and Fox (Conforti et al., 2020), and Lockdown in New York State (Miao et al., 2020). [NEED TO EDIT]
Appendix F BTSD Fine-tuning
Following TSE, we consider 4 datasets combined i.e. SemEval (Mohammad et al., 2016), AM (Stab et al., 2018), COVID-19 (Glandt et al., 2021), P-Stance (Li & Caragea, 2021), containing 19 targets, splitting samples (70-80%, 10-15%, 15-20%) depending on the dataset size into train, dev, and test respectively, maintaining the distribution from each of the targets. We will clarify it further in our revised version. [NEED TO EDIT]
Appendix G SemSim Details
Specifically, we employ the bert-base-uncased (Devlin et al., 2018) model to obtain contextual encoded representations of the target words. After performing a mean pooling operation on these representations, we apply cosine similarity (Singhal, 2001) to the encoded sequences to obtain the similarity score.
Appendix H Explanation with Samples in TG and SD
H.1 Target Generation
For instance, in the post "I mean, this reads to me like anyone who’s an alcoholic doesn’t go to heaven…", GPT-3.5 generated "Sinful behavior and heaven’s eligibility" as the target instead of the gold target "gay rights".
H.2 Stance Detection
For instance, in the explicit sample text, "Q. from audience: Malcom X said the most disrespected person in America is the black woman, has #MeToo changed this? Gender has been constructed in the image of the white woman, the idea of a level playing field, white women are often in better positions, says @HeidiMirza #CHevents", GPT-3.5 generated the target "Gender roles", which, while related, fails to capture the specific stance-related target "MeToo movement". This mistake reflects the model’s need for broader context to establish the correct link. Similarly, in the non-explicit sample "@DrOakley1689 @PhxApologetics Maybe they just misused their resources and their societies collapsed #deforestation #SemST", GPT-3.5 generated "Environmental degradation" as the target, while the gold target was "Climate change is a concern". Here, the generated target, though relevant as a subtopic, misses the broader context necessary to align with the gold target.
Appendix I Human Annotator Agreement
Dataset Description: The file includes a Tweet column with various Twitter posts. The Gold Target column represents the true target, indicating the primary topic or issue the tweet addresses, which in turn informs the stance noted in the Gold Stance column. For each tweet, there are eight generated targets (in columns T1 through T8) produced by different models, whose names remain anonymous for unbiased annotation. Additionally, there are empty cells (e.g., in the T1 Score column) designated for storing human evaluation scores corresponding to each generated target.
Definition of a Relevant Target: A generated target in the context of the target generation task is considered relevant if it accurately captures or closely aligns with the main topic, issue, or entity identified by the golden target in a given tweet. It should reflect the key subject matter or concern that the golden target represents, maintaining coherence with the overall context and meaning of the tweet.
Task Description: Your task is to assess the quality of each generated target by evaluating its relevance to the golden target in the tweet. You will do this by answering the following question and assigning a score to each generated target. Scoring criteria and examples are provided below.
“How closely does the generated target relate to the golden target in the tweet?”
Scoring Scale: 0 to 1 (Low to High)
  • 0 – Not Related: The generated target is entirely unrelated to the golden target. It does not reference or express anything connected to the intended topic.
  • 0.5 – Partially or Indirectly Related: The generated target has some relevance to the golden target but does not directly match the main concept. It may address a broader or narrower aspect of the topic, such as a parent topic, a subtopic, or a tangentially connected idea that is not fully aligned with the golden target’s core meaning.
  • 1 – Completely Related: The generated target is highly relevant to the golden target. It either uses a synonymous term, is semantically similar to the golden one, or conveys the same underlying idea, topic, or issue as the golden target. This means that the generated target captures the essence or main concept of the golden target, even if it uses different wording. For example, if the golden target is "climate change," a conceptually similar term could be "global warming" or "environmental crisis," as both refer to the broader issue of environmental concerns related to climate.
Example Scoring:
  1. 1.
    Tweet: Attempts to conceal the creationism-evolution controversy from students are dogmatic promotions of evolution. Not since blasphemy laws has competitive expression of thought been illegalized, and this is what evolutionists want to accomplish. This is evidenced by none other than the title of an evolutionist argument on this very page: "Schools should not teach theories that are completely at odds with each other",
    Gold Target: creationism,
    T1: creationism-evolution controversy, T1 Score: 1,
    T2: Education controversy, T2 Score: 0.5.
  2. 2.
    Tweet: If, on a supernatural level, lust of any sort counts the same consummated or unconsummated, why not just go ahead and consummate. You can’t get in any worse trouble, and if it makes you happy…,
    Gold Target: gay rights,
    T1: consummation, T1 Score: 0.5,
    T2: superstition, T2 Score: 0.
Appendix J Human Annotation Score Distribution
Figure 1: Distribution of human evaluation scores for target relevance assessment (0 – Not Related, 0.5 – Partially or Indirectly Related, 1 – Completely Related). The final score for each sample is determined by the majority vote among the three annotators. In cases of a tie, the average score of 0.5 is assigned. Complete details on human annotator agreement are provided in Appendix I.

### v3

@@
 Abstract
-Stance detection (SD) assesses a text’s position towards a target, typically labeled as "favor," "against," or "neutral." We introduce Open-Target Stance Detection (OTSD), where targets are neither seen during training nor provided as input.
-Evaluating Large Language Models (LLMs) like GPT-3.5, Llama 3, and Mistral, we compare their performance with the Target-Stance Extraction (TSE) approach, which has the advantage of using predefined targets.
-LLMs perform better than TSE in target generation when the real target is explicitly and not explicitly mentioned in the text.
-For stance detection, LLMs perform better in explicit scenarios but fail in non-explicit ones.
-[^1] We will release the code and the dataset.
+Stance detection (SD) identifies a text’s position towards a target, typically labeled as favor, against, or none.
+We introduce Open-Target Stance Detection (OTSD), the most realistic task where targets are neither seen during training nor provided as input.
+We evaluate Large Language Models (LLMs) GPT-4o, GPT-3.5, Llama-3, and Mistral, comparing their performance to the only existing work, Target-Stance Extraction (TSE), which benefits from predefined targets.
+Unlike TSE, OTSD removes the dependency of a predefined list, making target generation and evaluation more challenging.
+We also provide a metric for evaluating target quality that correlates well with human judgment.
+Our experiments reveal that LLMs outperform TSE in target generation when the real target is explicitly and not explicitly mentioned in the text.
+Likewise, for stance detection, LLMs excel in explicit cases with comparable performance in non-explicit in general.
+[^1] Dataset and code are available here: anonymous.4open.science/r/opentarget-5521
+[…]
+[…]
 Can Large Language Models Address Open-Target Stance Detection?
-Abu Ubaida Akash  and Ahmed Fahmy  and Amine Trabelsi {akaa2803, mosa2801, amine.trabelsi}@usherbrooke.ca Université de Sherbrooke
+Abu Ubaida Akash  Ahmed Fahmy  Amine Trabelsi Université de Sherbrooke {akaa2803,mosa2801,amine.trabelsi}@usherbrooke.ca
 1 Introduction
-Stance detection (SD) is the task of determining the position of a text or person towards a certain target, often split into “favor” and “against” with an optional third split being “neutral”, “none”, or “neither” (Sutter et al., 2024).
-The target word can be mentioned in the text explicitly, or sometimes the idea of the target word can be conveyed indirectly (Küçük and Can, 2020).
-In Zero-shot Stance Detection (ZSSD), the model predicts the stance for targets that it has not seen during the training phase.
-This task is essential because collecting training data for every possible topic or target is impractical (Allaway et al., 2021).
-While recent research has focused on ZSSD (Zhang et al., 2023; Li et al., 2023b; Wen and Hauptmann, 2023; Liang et al., 2022a; Zhu et al., 2022; Luo et al., 2022; Xu et al., 2022; Liang et al., 2022b; Li and Yuan, 2022), most studies assume that the target is known or manually identified and given as input, a situation that is uncommon in real-world applications.
-In practice, the target is unknown, and it is not necessarily mentioned explicitly in the text.
-Additionally, target annotation is a tedious and expensive task in SD (Küçük and Can, 2020; Sutter et al., 2024).
+Stance detection (SD) aims to determine the position of a text or person towards a certain target, typically categorized as “favor”, “against”, or “none”.
+The target can be mentioned explicitly in the text, or sometimes the idea of the target can be conveyed indirectly (Küçük and Can, 2020).
+In Zero-shot Stance Detection (ZSSD), a model predicts stances for targets it has not seen during training, which is crucial since collecting training data for every potential target is impractical (Allaway et al., 2021).
+While recent research has focused on ZSSD (Zhang et al., 2023; Li et al., 2023b; Wen and Hauptmann, 2023; Liang et al., 2022; Zhu et al., 2022; Luo et al., 2022; Xu et al., 2022), most studies assume that the target is known or manually identified and given as input, a rare scenario in real-world applications where the target is often unknown or not explicitly mentioned.
+Additionally, target annotation is an expensive task in SD (Küçük and Can, 2020).
 In this paper, we focus on a different yet challenging task we refer to as Open-Target Stance Detection (OTSD).
@@
 In OTSD, one important challenge lies in identifying the target from the text rather than having it provided as input.
-The target may be mentioned explicitly or not.
-The stance is predicted with respect to the produced target.
+The target may be mentioned explicitly or not and the stance is predicted with respect to the produced target.
 There has been one notable attempt to partially address this real-world setting of SD.
@@
 (2023a) generate targets from the input text and later detect stance based on these targets.
-However, before starting the SD task, they map these generated targets to a predefined list of golden targets (details in §2) to ensure exact wording and facilitate evaluation.
-They called their approach Target-Stance Extraction in the Wild (TSE).
-This approach does not fully align with our OTSD setting, which assumes no input target information is given during the whole process, and is impractical due to the need for a comprehensive list of all possible targets.
+However, in target generation, they map the generated targets to a predefined list of golden targets (details in §2) to ensure exact wording and facilitate evaluation.
+They call their approach TSE, which does not fully align with our OTSD setting.
+In OTSD we assume no input target information is given during the whole process, making it more practical than TSE, which requires a comprehensive list of all possible targets.
 Moreover, our experiment reveals a performance gap in Li et al.
-(2023a)’s approach when used on text with explicit or non-explicit target mentions.
-This aspect was not highlighted by the authors and warrants further investigation in the context of OTSD.
-In this work, we present and explore the OTSD task, focusing on its two main steps: Target Generation and Stance Detection (see Eq.
+(2023a)’s approach when used on text with explicit or non-explicit target mentions, a factor not addressed by the authors that warrants further investigation in the context of OTSD.
+In this work, we present and explore the OTSD task, focusing on its two main steps: Target Generation (TG) and Stance Detection (SD) (Eq.
 2).
 As a zero-shot task for both target and stance, Large Language Models (LLMs) appear well-suited for this challenge.
-Therefore, we examine the performance of LLMs and compare it to TSE, the primary existing work by Li et al.
-(2023a), which benefits from using a predefined list of targets during the process (see Eq.
+Therefore, we examine the performance of LLMs and compare it to the primary existing work, TSE, which benefits from using a predefined list of targets during the process (Eq.
 1).
-Our empirical study aims to address the following research questions: [RQ#1] How do proprietary and open LLMs—specifically GPT-3.5 (Brown et al., 2020), Llama 3 (Touvron et al., 2023), and Mistral (Jiang et al., 2023)—perform in open target generation compared to the TSE approach when the real target is explicitly or not explicitly mentioned in the text for stance classification, and how do they compare to each other?
-[RQ#2] How do the same LLMs perform compared to TSE in stance detection (in the context of OTSD) for both explicit and non-explicit cases, and how do they compare to each other?
-Our contributions consist of introducing the task of OTSD and conducting experiments and answer these research questions.
+Our empirical study aims to address the following research questions: [RQ#1] How do proprietary and open LLMs—specifically GPT-4o (Achiam et al., 2023), GPT-3.5 (Brown et al., 2020), Llama 3 (Touvron et al., 2023), and Mistral (Jiang et al., 2023)—perform in open TG compared to the TSE when the real target is explicitly or non-explicitly mentioned in the text for stance classification, and how do they compare to each other?
+[RQ#2] How do the same LLMs perform compared to TSE in SD (in the context of OTSD) for both explicit and non-explicit cases, and how do they compare to each other?
+Our contributions consist of introducing the task of OTSD, providing a target evaluation metric, and conducting experiments answering the research questions.
 2 Target-Stance Extraction (TSE)
@@
 For our comparison, we only consider their zero-shot setting.
-In TSE zero-shot, they first generate the target using keyphrase generation models, map the generated targets (multiple targets per sample) to a predefined list of golden targets to find the most similar one, and finally use it to detect the stance of the text toward the target (Eq.
+In TSE zero-shot, they first generate targets using keyphrase generation models, then map these targets to a predefined list of golden targets to find the closest match, and finally use this match to detect the stance of the text (Eq.
 1).
-The best match from their keyphrase model’s output, marked as TSE-BestGen, is used for our comparison.
-The best mapped target (TSE-mapped) is considered the predicted target.
-For example, given the text "Embracing different faiths teaches us empathy and respect.
-Let’s celebrate our diverse beliefs and foster unity.
-#Faith #Unity #Respect", TSE primarily generates targets such as Peace, Religious diversity, Respect, etc..
+The best match from the keyphrase model, marked as TSE-BestGen, and the best mapped target (TSE-mapped) is considered as the predicted target.
+For example, given the text "Embracing different faiths teaches … diverse beliefs and foster unity.", TSE primarily generates targets such as Peace, Religious diversity, Respect, etc..
 From a predefined list of possible targets (e.g., Face Mask, Atheism, Donald Trump, etc.), Atheism matches most closely with Religious diversity and is considered as the final predicted target.
-In OTSD, we aim at directly generating Religious diversity or a very semantically similar phrasing, avoiding the constraints of a predefined list.
-In the second step, they detect the stance toward their generated target.
+In OTSD, we aim to generate Religious diversity or a closely related phrase directly, bypassing the predefined list.
+The stance is then detected toward this generated target.
 3 Open-Target Stance Detection (OTSD)
@@
 To address the research questions outlined in §1, our approach to OTSD leverages the zero-shot learning capabilities of LLMs.
-We adopt the "Task Definition" prompting strategy for the OTSD task, as it has been shown by Cruickshank and Ng (2023) to outperform other strategies—such as "Task-only", "Context Analyze", "Context Question", "Zero-shot CoT", and "CoDA"—particularly for smaller models (7B-8B parameters) across most datasets.
-Given that this is an open-target task, we aim to use a straightforward prompting approach to maintain fairness and generalizability; we do not tailor the prompt for specific targets or include demonstrations or in-context examples, as this remains a zero-shot task.
-Since LLMs are sensitive to the length of generated text, we limit the target length to ensure a fair comparison with TSE in terms of target generation quality.
-To determine the optimal length, we test various target length instructions across all LLMs (§4.1) on 100 randomly selected samples from our dataset (§4.1), aiming to achieve an average length and standard deviation that closely match those of the golden targets.
-Based on our empirical study, we set the maximum target length to 5 for GPT-3.5 and 4 for Llama and Mistral to obtain an average length and standard deviations close to those of gold targets (explicit: $3.78\pm 1.5$; non-explicit 2.66±1.76)2.66$\pm$ 1.76)2.66 ± 1.76 ).
-Finally, we employ two different prompting approaches for target generation (TG) and stance detection (SD).
-TG+SD (Two-Step Approach)
-In this approach, we first generate the target and then sequentially detect the stance of the text toward the generated target.
+We adopt the "Task Definition" prompting strategy, as it has been shown by Cruickshank and Ng (2023) to outperform other strategies—such as "Task-only", "Context Analyze", "Context Question", "Zero-shot CoT", and "CoDA"—particularly for smaller models (7B-8B parameters) across most datasets.
+Since this is an open-target task, we use a straightforward prompting approach to ensure fairness and generalizability, avoiding tailored prompts for specific targets, or in-context examples, aligning with the zero-shot nature of this task.
+We employ two different prompting approaches for TG and SD.
+The prompts and their design justifications can be found in Appendix A.
+TG+SD (Two-Step Approach) In this approach, we first generate the target and then sequentially detect the stance of the text toward the generated target.
 By directing the LLM to focus on one task at a time, this method may reduce the cognitive load on the model.
-TG&SD (Single-Step Joint Approach)
-This approach combines target generation and stance detection into a single step using a unified prompt.
+TG&SD (Single-Step Joint Approach) This approach combines TG and SD into a single step using a unified prompt.
 By jointly generating the target and stance, the model can develop a more holistic understanding of the relationship between the text, target, and stance.
-The prompts and their design justifications can be found in Appendix A.
 4 Experiments
-In the OTSD task, we compare our approaches with TSE in both TG and SD.
+In OTSD task, we compare our approaches with TSE in both TG and SD.
 To ensure sensitivity and robustness, we run our experiments three times independently and report the average results  [^2] Score variations are negligible, up to four decimal places..
@@
 4.1 Dataset and Model
-We use the same dataset as TSE’s zero-shot setup for a fair comparison, consisting of a total of 3,000 samples considering three stance classes i.e., Favor, Against, and None.
-This includes 500 samples each from 6 different targets (see Appendix E).
-Additionally, we categorize the dataset into explicit and non-explicit samples.
-This categorization involves removing stop words and special characters from both texts and targets, lemmatizing all words, and subsequently comparing the text with the target.
-A sample is labeled as explicit if any target words are found in the corresponding text; otherwise, it is labeled as non-explicit.
-This process results in 1,804 explicit samples and 1,196 non-explicit samples (in Appendix E).
-It is important to note that because we compare our approach with TSE zero-shot, and they do not provided the code or sufficient details for reproducing the generation or selection of candidate targets, we are unable to use other stance detection datasets for benchmarking.
-In our focused contribution, we utilize both proprietary and non-proprietary LLMs, including GPT-3.5 (Brown et al., 2020), Llama-3 (Touvron et al., 2023), and Mistral (Jiang et al., 2023), to address our two research questions (§1).
-[^3] Although we do not claim that the results or findings are true for all LLMs, experimenting with other proprietary LLMs is financially constraining for us.
-Model descriptions can be found in Appendix  B.
+To ensure a fair comparison between TSE and LLM models, we use the same dataset as TSE’s zero-shot setup.
+VAST (Allaway and McKeown, 2020) is used to compare LLMs across different settings.
+Both TSE and VAST datasets contain three stance classes (Favor, Against, None), and contain 3,000 and 5,100 samples, respectively, in our study.
+VAST, originally a multi-target dataset, is converted to a single-target format (conversion process in Appendix E) to match the scope of our study.
+TSE has 500 samples for each of 6 targets (see Appendix F), while VAST includes 2,145 unique targets.
+We classify samples as explicit or non-explicit by removing stop words and special characters, lemmatizing all words, and checking for target words in the text.
+This results in 1,804 explicit and 1,196 non-explicit samples from TSE, and 3,120 explicit and 1,980 non-explicit samples from VAST (see Appendix F).
+Note that, due to the lack of code and details from TSE for reproducing target generation and selection, we cannot use other SD datasets for benchmarking.
+In our focused contribution, we use both proprietary and non-proprietary LLMs, including GPT-3.5, GPT-4o, Llama-3, and Mistral to address our two research questions (§1).
 4.2 Evaluation Methods
-To measure the generated target quality in OTSD task, we consider the following metrics:
-BTSD: We fine-tune the BERTweet model (Nguyen et al., 2020) following the same setup as TSE (fine-tuning details in Appendix F).
+BTSD: To measure the generated target quality, we fine-tune the BERTweet model (Nguyen et al., 2020) following the same setup as TSE (fine-tuning details in Appendix G).
 The SD score (F1-macro) is used as the evaluator for the generated targets.
 To assess the effectiveness of this metric, we experiment with different levels of target quality by modifying or removing words from the golden targets, substituting them with incorrect targets, and selecting random words from the vocabulary as targets.
-We find that the BERTweet F1 score consistently and significantly improved with higher-quality targets, in both explicit and non-explicit cases (Appendix C, Table 3).
-Therefore, the F1 score on stance detection serves as a reliable proxy for evaluating target generation quality in our dataset.
-Human Evaluation: To evaluate the quality of each generated target and to further validate the BTSD metric, we conduct a small-scale human evaluation using 500 randomly selected samples (300 explicit and 200 non-explicit, with 16.67% from each stance label).
-Three independent domain experts are asked to assess the relevance of the generated targets by evaluating its relevance to the golden target in the tweet.
-Each annotator is provided with an individual file, ensuring that their evaluations remain independent of one another.
-The full agreement details and score distribution are included in Appendix I and J, respectively.
+We find that the BERTweet F1 score consistently and significantly improved with higher-quality targets, in both explicit and non-explicit cases (see Appendix C).
+More importantly, BTSD score shows strong tau ($\tau$) correlation (KENDALL, 1938) of $0.85$ and $0.79$ for TSE and VAST, respectively, with human judgments gathered during experimentation (see the section below).
+Therefore, the BTSD serves as a reliable proxy for evaluating target generation quality.
+Human Evaluation: To evaluate the relevance of the generated targets and further validate the BTSD metric, we conduct a small-scale human evaluation using 500 randomly selected samples (300 explicit and 200 non-explicit) from both datasets, maintaining an equal ratio of the stance labels.
+Three annotators were asked to assess the relevance of the generated targets to the golden ones, classifying them as either 0 (Not Related), 0.5 (Partially Related), or 1 (Completely Related), following the guidelines provided in Appendix J.
+To measure inter-annotator agreement among the three raters, we compute Krippendorff’s $\alpha$ (Krippendorff, 2011), and Fleiss’s $\kappa$ (Fleiss, 1971) for each model, as shown in Appendix Table 5.
+We observe a solid overall agreement of $\alpha$=0.76 and $\kappa$=0.664.
 SemSim: In OTSD, since we do not rely on a predefined list of targets, the generated targets may not precisely match the gold targets but are semantically related.
 Therefore, we measure the semantic similarity between the generated and gold targets using BERT embeddings.
-Detailed steps can be found in Appendix G.
+Detailed steps can be found in Appendix H with human judgement correlation of $\tau$=$0.57$ for TSE, and $\tau$=$0.59$ for VAST.
 We use macro-average F1 for the task of SD.
-TG+SD TG&SD
-Model SemSim BTSD HEval SemSim BTSD HEval
-Explicit
-TSE-M 0.96 36.63 0.716 - - -
-TSE-B 0.86 35.8 0.338 - - -
-GPT-3.5 0.87 38.43 0.546 0.87 39.6 0.663
-Llama-3 0.88 38.31 0.54 0.87 37.75 0.558
-Mistral 0.85 38.47 0.586 0.87 39.02 0.65
-Non-explicit
-TSE-M 0.9 30.56 0.391 - - -
-TSE-B 0.82 29.32 0.215 - - -
-GPT-3.5 0.84 33.1 0.419 0.83 31.32 0.405
-Llama-3 0.85 33.06 0.447 0.83 33.81 0.49
-Mistral 0.83 30.18 0.386 0.83 34.12 0.495
-Table 1: Comparison of TG performance between TSE and LLMs across various evaluation metrics.
-HEval represents human evaluation.
-TSE-M and TSE-B correspond to TSE-mapped and TSE-BestGen, respectively.
-The highest score for each metric across all models is highlighted in bold, while for BTSD and HEval, the better results between the two approaches are underlined.
+TG+SD TG&SD TG+SD TG&SD
+Model SS BTSD HE SC SS BTSD HE SC SS BTSD HE SC SS BTSD HE SC
+Explicit Explicit
+[…]
+TSE-B 0.86 35.8 0.338 37.81 - - - - - - - - - - - -
+[…]
+[…]
+[…]
+[…]
+Non-explicit Non-explicit
+[…]
+TSE-B 0.82 29.32 0.215 31 - - - - - - - - - - - -
+[…]
+[…]
+[…]
+[…]
+Table 1: Comparison of TG and SD performance between TSE and LLMs using various metrics: TG is evaluated with SS (SemSim), BTSD (%), and HE (Human Evaluation) scores, while SC (%) measures SD performance.
+TSE-M and TSE-B represent TSE-mapped and TSE-BestGen, respectively.
+Metric-wise best results within a specific setting are underlined, and better results of a model across TG+SD and TG&SD for each dataset are in bold.
 5 Result Analysis
-5.1 Target Generation
-Following RQ#1, Table 1 shows that for the explicit case, Mistral achieves the highest BTSD score in the TG+SD setting, while GPT-3.5 leads in TG&SD.
-Overall, LLMs outperform TSE-BestGen in the BTSD metric and produce more relevant targets according to HEval scores.
-Although TSE-mapped has the highest SemSim and HEval scores, its BTSD score is lower due to generating either a perfect match or something entirely different (as illustrated in Figure 1).
-In contrast, the LLMs generate targets that are either highly related or partially related to the golden ones, with fewer irrelevant targets compared to TSE-mapped, leading to higher BTSD scores despite lower HEval.
-For non-explicit cases, LLMs outperform TSE-BestGen in both BTSD and HEval, with Llama-3 generating the most relevant targets (as indicated by the HEval score), even surpassing TSE-mapped.
-In the explicit case, TG&SD generally outperforms TG+SD across most LLMs, except for Llama, which performs better with TG+SD in BTSD.
-In non-explicit cases, GPT excels with TG+SD, while the other LLMs perform better with TG&SD.
-Among LLMs, GPT performs best overall in the explicit case, while Mistral leads in the non-explicit case.
-A manual inspection (on GPT-3.5’s 30 incorrect predictions) investigating the lower score in non-explicit cases reveals that the model struggles without sufficient context, often due to the way the TSE test set is constructed.
-This lack of context, such as implicit hints on the target or the context of other messages in the thread, causes the model to generate targets based solely on the provided text (example in Appendix H.1), leading to inaccuracies.
-Table 1 also confirms that higher HEval scores correlate with higher BTSD scores, validating BTSD as a reliable proxy for evaluating target quality.
-Model TG+SD TG&SD
-Explicit
-TSE-mapped 38.1 -
-TSE-BestGen 37.81 -
-GPT-3.5 42.68 47.61
-Llama-3 43.13 42.1
-Mistral 46.15 46.32
-Non-explicit
-TSE-mapped 32 -
-TSE-BestGen 31 -
-GPT-3.5 32.66 33.94
-Llama-3 33.36 31.9
-Mistral 33.61 30.47
-Table 2: Comparison of SD performance between TSE and LLMs, measured by F1-macro (in %).
-The highest scores across all models are highlighted in bold, with the best scores between the two approaches underlined.
-5.2 Stance Detection
-To address RQ#2, Table 2 shows that in stance detection, LLMs without a predefined list outperform TSE, which uses a predefined list, in both approaches (TG+SD and TG&SD) for explicit cases.
-In non-explicit cases, the scores are comparable in the TG+SD setting, but higher scores are observed for LLMs in the TG&SD setting.
-This is because, in explicit cases, LLMs generate better-quality targets than TSE-BestGen, leading to improved stance detection as evidenced in Table 1.
-For non-explicit cases, the holistic view provided by the TG&SD setting makes it easier for LLMs to infer stance from the text.
-Note that although LLM generated targets are better quality than TSE-BestGen, their stance labels can still be different from golden labels, resulting in low F1 scores.
-Our manual analysis revealed that these errors mainly arise from insufficient context in the input text, both explicit and non-explicit (see Appendix H.2 for examples).
-In both approaches, LLMs achieve much higher SD scores in explicit cases compared to non-explicit ones.
-Again, this is due to the incorrectly generated targets in the non-explicit case revealed by manual inspection.
-Incorrect targets correspond in a high percentage to incorrect stance labels.
-Most of targets were either completely incorrect or, in a few cases, contradictory to the real target, altering the stance.
+5.1 Target Generation (TG)
+As shown in Table 1, both human evaluation (HE) and BTSD indicate that LLMs generate higher-quality targets in explicit cases compared to TSE-BestGen, the best model for generating targets not matched to a predefined list.
+Although TSE-mapped achieves higher HE score in explicit case, its BTSD score is lower.
+This is because TSE-mapped produces either a perfect match (score=1) or a completely unrelated target (score=0) favouring higher HE score (see Appendix K).
+In contrast, LLMs mostly generate targets that are highly or partially related to golden targets, resulting in fewer irrelevant outputs and higher BTSD, despite slightly lower average HE scores.
+When the target is not explicitly mentioned, LLMs produce higher-quality targets than both TSE models.
+However, the quality drops for LLMs (TSE as well) compared to explicit cases (as per Table 1), with the best-performing LLM getting the HE score slightly lower than $0.5$ (refers to partially related target) on average.
+A manual inspection of GPT-3.5’s 30 incorrect predictions suggests that the lower score in non-explicit cases is due to insufficient context, often resulting from how the TSE test set is constructed.
+The lack of implicit hints or surrounding context leads the model to generate targets based solely on the text provided, causing inaccuracies (see example in Appendix I.1).
+When comparing tested proprietary and open LLMs using BTSD and HE, there is no clear advantage in TG (i.e., not all proprietary models outperform open ones, or vice versa).
+We note that GPT-4o consistently produces targets of equal or higher quality than other tested models across the settings, including explicit and non-explicit target mentions, TG+SD, and TG&SD approaches, and across the datasets.
+Among the tested open LLMs, Mistral generates higher quality targets than Llama-3 in most settings.
+Comparing between the approaches, most models perform better in TG&SD across both datasets, for explicit and non-explicit cases.
+5.2 Stance Detection (SD)
+As shown in Table 1, according to the SC (stance classification), LLMs outperform both TSE models in detecting stances for explicit cases, while performing comparably with TSE models in non-explicit cases (except GPT-4o).
+In general, LLMs perform better in SD when the target is explicitly mentioned in the text compared to when it is not.
+Our manual analysis shows that in non-explicit cases, the input text lacks sufficient context about the target, leading to lower SC scores.
+In general, the experimented proprietary models (GPT-3.5 and GPT-4o) outperform the open ones in TG&SD in both the explicit and non-explicit cases across the datasets.
+In some cases of TG+SD setting, though, the open LLMs achieve comparatively higher scores.
+Comparing approaches, LLMs tend to perform better when prompted with the holistic TG&SD, though their TG+SD results remain competitive, specifically in non-explicit cases across both datasets.
+Although the quality of the generated targets correlates positively with the SC score in explicit cases (TSE $\tau$=$0.71$, VAST $\tau$=$0.29$), this correlation is negative in non-explicit cases (TSE $\tau$=$-0.14$, VAST $\tau$=$-0.29$).
+We manually investigate the negative correlation (lower SC with higher quality target) in non-explicit cases.
+Indeed, LLMs such GPT-4o may generate highly relevant targets that are closely related to the golden targets but nearly ‘antithetical’ or opposite, leading to a stance reversal in comparison to the golden stance.
+For example, when the golden target is "permits to carry guns", GPT-4o generates "gun control in universities", while GPT-3.5 generates "guns".
+Although GPT-4o’s target is more closely related to the golden target, it represents the opposite stance.
+On the other hand, GPT-3.5’s target, though broader, less accurate and less related, aligns with the stance of golden target.
+This highlights the need for a coherence measure in OTSD that accounts for the target-stance alignment.
 6 Conclusion
-LLMs outperform TSE in target generation for both explicit and non-explicit cases, despite TSE’s use of a predefined list.
-However, LLMs struggle more with generating targets when they are not explicitly mentioned in the text.
-LLMs also perform better in stance detection than TSE for both explicit and non-explicit cases.
-Additionally, LLMs show better stance detection performance when the target is explicitly mentioned compared to non-explicit cases.
-Finally, the single-step joint prompt approach (TG&SD) is more effective than the two-step approach (TG+SD) for both TG and SD in both explicit and non-explicit scenarios.
+We introduce the Open Target Stance Detection (OTSD) task, where targets are neither seen during training nor provided as input, making it more realistic and challenging.
+We examine proprietary and open LLMs in our OTSD task, and propose a metric to evaluate the quality of Target Generation (TG).
+We compare to TSE, a method leveraging a predefined list of targets.
+LLMs outperform it in TG.
+However, LLMs struggle in generating quality targets when these are not explicitly mentioned.
+LLMs also surpass TSE in Stance Detection (SD) when the target is explicitly mentioned while performing comparably in non-explicit cases.
+Finally, a single-step joint prompt approach (TG&SD) of LLMs proves more effective than a two-step approach (TG+SD) for both TG and SD.
 Limitation
-We do not provide a specific coherence measure to evaluate the Open-Target Stance Detection task, as the target and the stance toward it are inherently interconnected.
-Due to resource constraints, our experiments were conducted on a small scale in terms of the models used.
-The human evaluation of target quality was also limited to 500 samples.
+The limitations of our work can be listed as follows:
+  • We did not provide a specific coherence measure to evaluate the Open-Target Stance Detection task, which is worth to be investigated in the future work, as the target and the stance toward it are inherently interconnected.
   • Regarding potential data leakage during the pre-training of LLMs, it is possible that these models were exposed to some of our test data, leading to prior knowledge of certain entities.
 However, since the LLMs were not explicitly pre-trained on the same task, we continue to refer to our approach as zero-shot stance detection.
+  • Although we did not claim that the results or findings are true for all LLMs, experimenting with other proprietary LLMs was financially constraining for us.
 Appendix A Prompting Details
+Since LLMs are sensitive to the length of generated text, we limit the target length to ensure a fair comparison with TSE in terms of target generation quality.
+To determine the optimal length, we test various target length instructions across all LLMs (§4.1) on 100 randomly selected samples from our dataset (§4.1), aiming to achieve an average length and standard deviation that closely match those of the golden targets.
+Based on our empirical study, we set the maximum target length to 5 for GPT-3.5 and 4 for Llama and Mistral to obtain an average length and standard deviations close to those of gold targets (explicit: $3.78\pm 1.5$; non-explicit 2.66±1.76)2.66$\pm$ 1.76)2.66 ± 1.76 ).
 TG+SD (Two-Step Approach)
@@
 This model is decoder-only, similar to models like GPT-3 (Brown et al., 2020) and GPT-4 (Achiam et al., 2023), but with optimizations for efficiency and performance.
+GPT-4o
+is a decoder-only model in the GPT series with an optimized architecture designed to enhance efficiency and accuracy.
+It builds upon the foundations of GPT-4 (Achiam et al., 2023), incorporating improvements in training and fine-tuning for better performance across diverse language tasks.
 Llama-3-8B Instruct
@@
 Appendix C BTSD Metric Justification
-Input F1 (macro)
+Input F1-macro (in %)
 Explicit Non-explicit
 Tweet only 31.02 30.37
--w/ Gold Target 41.67 34.65
+-w/ Gold Target (GT) 41.67 34.65
 -w/ altered GT 35.3 32.54
@@
 -w/ random vocab 18.66 14
-Table 3: Macro avg. F1-Score of BertTweet on Stance Detection with different quality levels of the input target.
+Table 2: Macro avg. F1 score of BertTweet in detecting stances with different quality levels of the input target.
 Table 2 contains the results of the stratified experiments with different measure stance detection F1 score when giving as input the text and a target, varying the quality of this latter.
@@
 All of the experiments were run on a computer with Ubuntu 22.04 Linux with x64 CPU with 4 cores, 32 GB of RAM and one NVIDIA V100 GPU.
-Appendix E Dataset Samples
-Targets: Creationism (Somasundaran and Wiebe, 2010), Gay Rights (Somasundaran and Wiebe, 2010), Climate Change is a Concern (Mohammad et al., 2016), MeToo Movement (Gautam et al., 2020), Merger of Disney and Fox (Conforti et al., 2020), and Lockdown in New York State (Miao et al., 2020).
-[NEED TO EDIT]
-Appendix F BTSD Fine-tuning
+Appendix E VAST Dataset Processing
+Originally, the VAST dataset is a multi-target dataset, with multiple targets assigned to each sample text.
+However, since our work focuses on single-target analysis, we processed the VAST dataset accordingly to align it with our study (see Table 4).
+In VAST, there are two types of target columns: "ori_topic" which contains targets heuristically extracted, and "new_topic," which contains the final annotated targets determined by annotators.
+The "ori_topic" often includes duplicate targets for each text, while the "new_topic" provides unique targets with different stances.
+To convert the dataset to a single-target format, we first filtered the samples to retain only the most frequent stance within each group of duplicate "ori_topic" entries.
+Next, we computed the average semantic similarity of each "new_topic" within these duplicate groups using BERT embeddings.
+For groups containing only a single row, the sample was directly retained.
+Then, we calculated the average pairwise cosine similarity for each "new_topic." Finally, within each duplicate "ori_topic" group, we kept only the row with the highest average semantic similarity score.
+After this conversion, we obtained a total of 5,100 samples (as shown in Appendix Table 4), covering both explicit and non-explicit samples (the filtering process for explicit and non-explicit samples is described in Section 4.1) from the original 18,545 samples.
+Appendix F Dataset Samples
+Following TSE, our targets are Creationism (Somasundaran and Wiebe, 2010), Gay Rights (Somasundaran and Wiebe, 2010), Climate Change is a Concern (Mohammad et al., 2016a), MeToo Movement (Gautam et al., 2020), Merger of Disney and Fox (Conforti et al., 2020), and Lockdown in New York State (Miao et al., 2020).
+Samples from TSE and VAST in both explicit and non-explicit settings are given in Table 3 and 4, respectively.
+Target Text Stance
+Explicit
+creationism
+It is not ’appropriate’ to teach creationism as a means for upholding the bible. This pre-supposes that the
+bible should be upheld in a literal sense. Many Christians object to this. More importantly, governments
+should not be in the business of upholding the Bible.
+AGAINST
+merger of disney and fox
+@ComicBookNOW: FOX reportedly wants to make that deal with DISNEY! X-MEN in the MCU is
+now closer than ever
+FAVOR
+gay rights
+Most health care organizations support gay parenting as equally capable
+as heterosexual parenting These organizations are the most capable of determining the capabilities of
+homosexuals to perform dutifully as parents.
+FAVOR
+lockdown in new york state
+@adamajacoby @RV1026 @CNNPolitics Apparently, better than the "news sources" that keep you
+informed. You didn’t even know America had to lock down. You didn’t even know Cuomo put
+COVID patients in nursing homes. They are keeping you misinformed and looking like a
+complete moron. Do better!
+FAVOR
+metoo movement
+@KRKBoxOffice Why everyone was silent when she molested Saif Sir ? Why no #MeToo campaign
+that time??
+AGAINST
+climate change is a concern
+Being an engaged mom, means voting 4 the climate 2. Supporting only candidates who have a plan 2 act
+on #playin4climate #SemST
+FAVOR
+Non-explicit
+creationism
+My point was not that Genesis contradicted itself but that if you took it literally it contradicted itself. The
+problem is not with Genesis so much as a literal interpretation of Genesis. You have to get away from
+literalism to make sense of the two creation stories, with their different orders of creation. You have
+proved my point by showing how a less than literal interpretation of the passages gets round the problem
+of a literal interpretation.
+AGAINST
+gay rights
+That’s fine. I support that.Here in California, we have a prospective bill that would do just that.I honestly
+think marriage is a religious thing, but religion is in the eye of the beholder.
+FAVOR
+lockdown in new york state
+RT @Lukewearechange: So NYC announced another lock down coming this Sunday evening, I was
+suppose to be getting out of here with a friend
+FAVOR
+climate change is a concern Considering moving yo Antarctica as thats the only way I could possibly become more #chill #SemST NONE
+Table 3: Samples from TSE dataset with all the 6 targets in both explicit and non-explicit settings.
+Target Text Stance
+Explicit
+health care law
+Congress should delay the law for a year…"? Tell that to my niece whose 17 month old toddler has
+just been diagnosed with a rare disease that will need years of medical care. Without the provisions
+in the new health care law, this family would be facing bankruptcy. I prefer living in a nation where
+we have the decency to realize that health care is a right. Obviously those who seem to be objecting
+to the new health care law have never faced what 30 million people face in our country every day.
+For them another year’s delay is life or death.
+FAVOR
+facebook
+"…Facebook gets a bad rap; it didn’t cause the cheating. It just made it more convenient to do
+(and perhaps easier to catch)." Hmm, call me crazy, but Facebook shares some of the responsibility
+– does it not? It’s like refusing to assign blame to gun dealers and drug dealers. They increase
+accessibility to illicit materials, which essentially furthers the end goal of usage. Now, this isn’t to
+suggest that Facebook is entirely at fault, but it does play an active role.
+AGAINST
+stability of the economy
+"…one must ask how much money they must make to demonstrate that they are among the best
+managed companies on the planet." They must make enough money to insure that they can never
+fail and threaten the stability of the worlds economy again. That much money.
+NONE
+palestinian authority
+"But let’s start with the basics. Any nation wishing to declare independence should meet three
+essential elements: a strong central government, control of defined territory and security. The
+Palestinian Authority does not yet meet any of them." Historians would strongly disagree with you.
+Around sometime late eighteenth century some thirteen odd British colonies did "declare
+independence" without any resemblance of "strong central government" with its territory controlled
+by the British. Today a certain French foreign minister called it a "hyper-power" and not a
+"failed state". Please check upon a historian or read up your history. Trust me! It is true!!
+AGAINST
+economists
+"Economists do certainly over-reach sometimes. We tend to apply the lens of economic efficiency to
+situations where many people apply the lens of fairness." I get a kick out of economists believing
+that real live humans are rational economic actors. Really?
+AGAINST
+Non-explicit
+restaurant
+"…tipping motivates people who work long, busy hours catering to the needs of others. It’s the best way
+to ensure optimal service…" By this logic anyone who works long, busy hours catering to the needs of
+others should be tipped. Tip the doctor. Tip the grocery clerk. Tip the airline counter agent. Tip the
+airline pilot. Etc. I fail to see why those whose particular service happens to be delivering plates of
+food warrant their own method of compensation. One that puts an onus of extra calculation and
+deliberation on every single customer, every time they sit down to eat and relax. Thankfully, most other
+services in this world are one-price to the customer. It’s left to the employer to do the work of assessing
+whether the employee is providing good service.
+AGAINST
+prostitution
+"Granted, legalizing the profession might make it attractive for sex traffickers but the benefits outweigh
+this prospect." !!? The benefits of receiving tax revenue for the state outweighs the negative aspects of
+having 12 year old girls being sex trafficked into brothels and coerced to work on the streets?! What
+kind of logic is that, and what kid of person are you Ms. Unigwe?
+AGAINST
+vaccination
+"It is a news media-driven misperception that parents who claim philosophical or religious exemptions
+are uneducated or misinformed. MOST PARENTS WHO INDIVIDUALIZE THE VACCINE
+SCHEDULE ARE ACTIVELY EDUCATING THEMSELVES, CONTINUALLY ASSESSING THEIR
+FAMILY’S SPECIFIC HEALTH NEEDS, and doing everything they can to keep their children safe and
+healthy." Ms. Margulis offers no data to support her blanket assertion about the industry and motivation
+of "most parents," a position which, on its face, seems improbable given the breadth and weight of
+scientific evidence supporting immunization which is not subject to reasonable dispute. Based on anecdotal
+information, "most parents" are refusing to vaccinate their children based on their gross misunderstandings
+and unwarranted fears of the alleged risks of immunization, or based on their unique interpretations of
+religious dictates. If you intend to offer relevant commentary, Ms. Margulis, it should be evidence-based.
+Your personal opinions are no more interesting (or informed) than mine.
+FAVOR
+mentally
+’…food pornography, musical pornography, mental pornography…’ And yet I have never been in a public
+library and seen a man at a computer masturbating under his coat to videos of food, music, or anything
+’mental’, so let’s no pretend they are all the same beast, okay?
+NONE
+Table 4: Samples from VAST dataset with few different targets in both explicit and non-explicit settings.
+Appendix G BTSD Fine-tuning
 Following TSE, we consider 4 datasets combined i.e.
-SemEval (Mohammad et al., 2016), AM (Stab et al., 2018), COVID-19 (Glandt et al., 2021), P-Stance (Li & Caragea, 2021), containing 19 targets, splitting samples (70-80%, 10-15%, 15-20%) depending on the dataset size into train, dev, and test respectively, maintaining the distribution from each of the targets.
-We will clarify it further in our revised version.
-[NEED TO EDIT]
-Appendix G SemSim Details
+SemEval (Mohammad et al., 2016b), AM (Stab et al., 2018), COVID-19 (Glandt et al., 2021), P-Stance (Li and Caragea, 2021), containing 19 targets, splitting samples by (70%-train, 10%-dev, 20%-test), maintaining equal distribution from each of the targets.
+To assess the quality of the generated targets, we train a stance classifier model (referred as BTSD) as follows.
+Given a text $x$ and a target $t$, we first formulate the input as a sequence $s=[[CLS]t[SEP]x]$ where $[CLS]$ is a token that encodes the sentence and $[SEP]$ is used to separate the sentence $x$ and the target $t$.
+Then the representation of $[CLS]$ token is used to predict the stance toward target $t$.
+Note that $t$ is the golden target here.
+Appendix H SemSim Details
 Specifically, we employ the bert-base-uncased (Devlin et al., 2018) model to obtain contextual encoded representations of the target words.
 After performing a mean pooling operation on these representations, we apply cosine similarity (Singhal, 2001) to the encoded sequences to obtain the similarity score.
-Appendix H Explanation with Samples in TG and SD
-H.1 Target Generation
+Appendix I Explanation with Samples in TG and SD
+I.1 Target Generation
 For instance, in the post "I mean, this reads to me like anyone who’s an alcoholic doesn’t go to heaven…", GPT-3.5 generated "Sinful behavior and heaven’s eligibility" as the target instead of the gold target "gay rights".
-H.2 Stance Detection
+I.2 Stance Detection
 For instance, in the explicit sample text, "Q.
@@
 This mistake reflects the model’s need for broader context to establish the correct link.
-Similarly, in the non-explicit sample "@DrOakley1689 @PhxApologetics Maybe they just misused their resources and their societies collapsed #deforestation #SemST", GPT-3.5 generated "Environmental degradation" as the target, while the gold target was "Climate change is a concern".
-Here, the generated target, though relevant as a subtopic, misses the broader context necessary to align with the gold target.
-Appendix I Human Annotator Agreement
+Appendix J Human Annotator Agreement
+TSE VAST
+Model KA FK KA FK KA FK KA FK
+TSE-M (TG+SD) 1 1 1 1 - - - -
+TSE-B (TG+SD) 0.749 0.633 0.855 0.774 - - - -
+GPT-3.5 (TG+SD) 0.726 0.603 0.882 0.722 0.788 0.651 0.762 0.649
+GPT-3.5 (TG&SD) 0.82 0.783 0.709 0.616 0.819 0.72 0.714 0.626
+GPT-4 (TG+SD) 0.746 0.63 0.717 0.611 0.79 0.668 0.785 0.66
+GPT-4 (TG&SD) 0.857 0.75 0.727 0.631 0.852 0.785 0.794 0.681
+Lama-3 (TG+SD) 0.823 0.723 0.841 0.762 0.736 0.634 0.764 0.65
+Lama-3 (TG&SD) 0.827 0.758 0.708 0.617 0.748 0.639 0.751 0.647
+Mistral (TG+SD) 0.884 0.762 0.768 0.709 0.786 0.661 0.724 0.635
+Mistral (TG&SD) Explicit 0.836 0.786 Non-explicit 0.732 0.687 Explicit 0.81 0.707 Non-explicit 0.796 0.694
+Table 5: Inter-annotator reliability score based on Krippendorff’s $\alpha$ (KA) and Fleiss’s $\kappa$ (FK) coefficient metrics across all the combinations in two datasets.
+TSE-mapped and TSE-BestGen are reffered by TSE-M and TSE-B, respectively.
+The reliability score is 1 for TSE-M as it classifies between the golden targets instead of generating.
 Dataset Description: The file includes a Tweet column with various Twitter posts.
@@
     T2: superstition, T2 Score: 0.
-Appendix J Human Annotation Score Distribution
+Appendix K Human Annotation Score Distribution
 Figure 1: Distribution of human evaluation scores for target relevance assessment (0 – Not Related, 0.5 – Partially or Indirectly Related, 1 – Completely Related).
@@
 In cases of a tie, the average score of 0.5 is assigned.
-Complete details on human annotator agreement are provided in Appendix I.
+The details on human annotator guideline are provided in Appendix J, and the annotator agreement is in Appendix Table 5.

### v4

@@
 Can Large Language Models Address Open-Target Stance Detection?
-Abu Ubaida Akash  Ahmed Fahmy  Amine Trabelsi Université de Sherbrooke {akaa2803,mosa2801,amine.trabelsi}@usherbrooke.ca
+Abu Ubaida Akash  Ahmed Fahmy  Amine Trabelsi Department of Computer Science, Université de Sherbrooke {akaa2803,mosa2801,amine.trabelsi}@usherbrooke.ca
 1 Introduction
@@
 Human Evaluation: To evaluate the relevance of the generated targets and further validate the BTSD metric, we conduct a small-scale human evaluation using 500 randomly selected samples (300 explicit and 200 non-explicit) from both datasets, maintaining an equal ratio of the stance labels.
-Three annotators were asked to assess the relevance of the generated targets to the golden ones, classifying them as either 0 (Not Related), 0.5 (Partially Related), or 1 (Completely Related), following the guidelines provided in Appendix J.
+Three annotators were asked to assess the relevance of the generated targets to the golden ones, classifying them as either 0 (Not Related), 0.5 (Partially Related), or 1 (Completely Related), following the guidelines provided in Appendix K.
 To measure inter-annotator agreement among the three raters, we compute Krippendorff’s $\alpha$ (Krippendorff, 2011), and Fleiss’s $\kappa$ (Fleiss, 1971) for each model, as shown in Appendix Table 6.
@@
 Although TSE-mapped achieves higher HE score in explicit case, its BTSD score is lower.
-This is because TSE-mapped produces either a perfect match (score=1) or a completely unrelated target (score=0) favouring higher HE score (see Appendix K).
+This is because TSE-mapped produces either a perfect match (score=1) or a completely unrelated target (score=0) favouring higher HE score (see Appendix M).
 In contrast, LLMs mostly generate targets that are highly or partially related to golden targets, resulting in fewer irrelevant outputs and higher BTSD, despite slightly lower average HE scores.
@@
 A manual inspection of GPT-3.5’s 30 incorrect predictions suggests that the lower score in non-explicit cases is due to insufficient context, often resulting from how the TSE test set is constructed.
-The lack of implicit hints or surrounding context leads the model to generate targets based solely on the text provided, causing inaccuracies (see example in Appendix I.1).
+The lack of implicit hints or surrounding context leads the model to generate targets based solely on the text provided, causing inaccuracies (see example in Appendix J.1).
 When comparing tested proprietary and open LLMs using BTSD and HE, there is no clear advantage in TG (i.e., not all proprietary models outperform open ones, or vice versa).
@@
 Our manual analysis shows that in non-explicit cases, the input text lacks sufficient context about the target, leading to lower SC scores.
-In general, the experimented proprietary models (GPT-3.5 and GPT-4o) outperform the open ones in TG&SD in both the explicit and non-explicit cases across the datasets.
+Overall, the experimented proprietary models (GPT-3.5 and GPT-4o) outperform the open ones in TG&SD in both the explicit and non-explicit cases across the datasets.
 In some cases of TG+SD setting, though, the open LLMs achieve comparatively higher scores.
@@
   • We did not provide a specific coherence measure to evaluate the Open-Target Stance Detection task, which is worth to be investigated in the future work, as the target and the stance toward it are inherently interconnected.
+  • We conducted a thorough investigation of simple prompting for LLMs, inspired from the stance detection literature (§3.2), which has been shown to outperform other strategies, and provides the answers of our research questions already.
+However, there might be still room for improvement through a more detailed examination of different prompting techniques or task-specific pre-training.
+While we present some preliminary results on the chain-of-thought approach in Appendix L, further investigation is needed to fully understand its potential.
   • Regarding potential data leakage during the pre-training of LLMs, it is possible that these models were exposed to some of our test data, leading to prior knowledge of certain entities.
@@
 Originally, the VAST dataset is a multi-target dataset, with multiple targets assigned to each sample text.
-However, since our work focuses on single-target analysis, we processed the VAST dataset accordingly to align it with our study (see Table 4).
+However, since our work focuses on single-target analysis, we processed the VAST dataset accordingly to align it with our study (samples in Table 4).
 In VAST, there are two types of target columns: "ori_topic" which contains targets heuristically extracted, and "new_topic," which contains the final annotated targets determined by annotators.
@@
 After performing a mean pooling operation on these representations, we apply cosine similarity (Singhal, 2001) to the encoded sequences to obtain the similarity score.
-Appendix I Explanation with Samples in TG and SD
-I.1 Target Generation
+Appendix I OTSD with EZSTANCE
+We explore the OTSD task using EZSTANCE (Zhao and Caragea, 2024), the largest zero-shot stance detection (ZSSD) dataset, to address how proprietary and open LLMs compare in both target generation (TG) and stance detection (SD) tasks when the target is explicitly or implicitly mentioned in the text.
+Given the open-target nature of our task, EZSTANCE was selected for its diverse range of domains and unique target set, reflecting real-world scenarios.
+I.1 Dataset Processing
+The dataset originally contains 47,316 Twitter samples across three stance categories (favor, against, and neutral).
+Since EZSTANCE features multiple targets for some texts, we adapted it into a single-target dataset (one target per text) to better align with our research by following a series of steps.
+The original dataset is organized into two subtasks (target-based and domain-based), with each subtask containing two target types: claim and noun-phrase.
+First, we combine the samples from both target types and remove duplicates.
+For instances where both target types exist, we retain the noun-phrase version, as it poses a greater challenge in generating precise targets.
+Then, we merge the samples from both subtasks and deduplicate them based on their first occurrence.
+This process results in 9,462 unique samples with 6,873 distinct targets.
+We further divide the dataset into explicit and non-explicit cases, as outlined in Section 4.1, yielding 9,313 explicit and 149 non-explicit samples, where the target is mentioned explicitly or implicitly in the text, respectively.
+I.2 Result Analysis
+As shown in Table 5, LLMs face challenges in target generation for non-explicit samples compared to when the target is explicitly mentioned, based on BTSD and HE scores.
+When comparing proprietary and open models, consistent with our earlier findings (§5.1), no single LLM consistently outperforms the others.
+While GPT-4o tends to generate higher-quality targets in the TG+SD approach, Mistral surpasses other models in some TG&SD cases.
+Across both explicit and non-explicit samples, all models perform noticeably better using the TG&SD approach.
+For stance detection, as shown in 5, similar to our findings with TSE and VAST in Section 5.2, LLMs perform better in SC when the target is explicitly mentioned.
+When comparing proprietary and open models, GPT-4o and Mistral both demonstrate competitive performance across approaches.
+Overall, LLMs consistently achieve better stance detection in the TG&SD approach.
+TG+SD TG&SD
+Model SS BTSD HE SC SS BTSD HE SC
+Explicit
+[…]
+[…]
+[…]
+[…]
+Non-explicit
+GPT-3.5 0.76 38.78 0.377 36.69 0.81 41.89 0.402 38.97
+[…]
+[…]
+[…]
+Table 5: Comparison of TG and SD performance on EZSTANCE dataset (Zhao and Caragea, 2024) among the LLMs using various metrics: TG is evaluated with SS (SemSim), BTSD (%), and HE (Human Evaluation) scores, while SC (%) measures SD performance.
+Metric-wise best results within a specific setting are underlined, and better results of a model across TG+SD and TG&SD for each dataset are in bold.
+Appendix J Explanation with Samples in TG and SD
+J.1 Target Generation
 For instance, in the post "I mean, this reads to me like anyone who’s an alcoholic doesn’t go to heaven…", GPT-3.5 generated "Sinful behavior and heaven’s eligibility" as the target instead of the gold target "gay rights".
-I.2 Stance Detection
+J.2 Stance Detection
 For instance, in the explicit sample text, "Q.
@@
 This mistake reflects the model’s need for broader context to establish the correct link.
-Appendix J Human Annotator Agreement
+Appendix K Human Annotator Agreement
 TSE VAST
@@
     T2: superstition, T2 Score: 0.
-Appendix K Human Annotation Score Distribution
+Metrics TG+SD TG&SD
+Ex N-Ex Ex N-Ex
+BTSD-TSE 38.47 30.73 37.88 32.14
+BTSD-VAST 41.10 41.54 41.34 40.38
+SC-TSE 47.36 38.11 49.14 36.00
+SC-VAST 48.47 46.46 49.00 45.41
+Table 7: Caption goes here TG (evaluated by BTSD) and SD (evaluated by SC) performance using COT prompting on both TSE and VAST datasets in both Explicit and Non-explicit cases across the two approaches: TG+SD, TG&SD.
+Ex and N-Ex represent Explicit and Non-explicit cases, respectively.
+Appendix L COT Initial Result
+We initiate an exploration of alternative prompting techniques by employing the chain-of-thought (COT) prompting strategy for both target generation (TG) and stance detection (SD) tasks, across two approaches: TG+SD and TG&SD.
+As an initial phase, in Table 7, we assess target quality using the BTSD metric and evaluate stance detection performance with the SC metric, applied to both the TSE and VAST datasets for explicit and non-explicit cases.
+Appendix M Human Annotation Score Distribution
+After conducting the human annotation for target quality assessment, we plot the score distribution provided by the annotators, as shown in Figure 1.
+We observe that LLMs generally produce targets that are either partially or fully related to the golden targets, outperforming the TSE models in this regard.
+The targets generated by TSE-mapped are either identical to the golden targets or entirely unrelated, as TSE-mapped selects targets in its final stage (as described in Section 2) from a pre-defined list.
+In contrast, the targets generated by TSE-BestGen are often either partially or completely unrelated to the golden targets, indicating lower quality compared to LLMs.
 Figure 1: Distribution of human evaluation scores for target relevance assessment (0 – Not Related, 0.5 – Partially or Indirectly Related, 1 – Completely Related).
@@
 In cases of a tie, the average score of 0.5 is assigned.
-The details on human annotator guideline are provided in Appendix J, and the annotator agreement is in Appendix Table 5.
+The details on human annotator guideline are provided in Appendix K, and the annotator agreement is in Appendix Table 6.
+Appendix N Vsiual Comparison of TG and SD Performance
+To better illustrate the data in Table 1, we have plotted Figures 2 and 3, which represent the TG and SD performance, respectively.
+These figures provide a clearer visualization of the conclusions discussed in Sections 5.1 and 5.2.
+Figure 2: Comparison of TG performance of TSE and LLMs, as well as the performance among LLMs across both the TSE and VAST datasets, using various metrics: SS (Semantic Similarity), BTSD (%), and HE (Human Evaluation).
+The graphs are generated based on the data in Table 1.
+Figure 3: Comparison of SD performance of TSE and LLMs, as well as the performance among LLMs across both the TSE and VAST datasets, using SC (stance classification) score.
+The graphs are generated based on the data in Table 1.

### v5

@@
 We introduce Open-Target Stance Detection (OTSD), the most realistic task where targets are neither seen during training nor provided as input.
-We evaluate Large Language Models (LLMs) GPT-4o, GPT-3.5, Llama-3, and Mistral, comparing their performance to the only existing work, Target-Stance Extraction (TSE), which benefits from predefined targets.
+We evaluate Large Language Models (LLMs) from GPT, Gemini, Llama, and Mistral families, comparing their performance to the only existing work, Target-Stance Extraction (TSE), which benefits from predefined targets.
 Unlike TSE, OTSD removes the dependency of a predefined list, making target generation and evaluation more challenging.
 We also provide a metric for evaluating target quality that correlates well with human judgment.
-Our experiments reveal that LLMs outperform TSE in target generation when the real target is explicitly and not explicitly mentioned in the text.
-Likewise, for stance detection, LLMs excel in explicit cases with comparable performance in non-explicit in general.
-[^1] Dataset and code are available here: anonymous.4open.science/r/opentarget-5521
+Our experiments reveal that LLMs outperform TSE in target generation, both when the real target is explicitly and not explicitly mentioned in the text.
+Similarly, LLMs overall surpass TSE in stance detection for both explicit and non-explicit cases.
+However, LLMs struggle in both target generation and stance detection when the target is not explicit.
+[^1] Dataset and code are available here: anonymous.4open.science/r/opentarget-5521.
 […]
@@
 In Zero-shot Stance Detection (ZSSD), a model predicts stances for targets it has not seen during training, which is crucial since collecting training data for every potential target is impractical (Allaway et al., 2021).
-While recent research has focused on ZSSD (Zhang et al., 2023; Li et al., 2023b; Wen and Hauptmann, 2023; Liang et al., 2022; Zhu et al., 2022; Luo et al., 2022; Xu et al., 2022), most studies assume that the target is known or manually identified and given as input, a rare scenario in real-world applications where the target is often unknown or not explicitly mentioned.
-Additionally, target annotation is an expensive task in SD (Küçük and Can, 2020).
+While recent research has focused on ZSSD (Zhang et al., 2023; Li et al., 2023b; Wen and Hauptmann, 2023; Liang et al., 2022; Zhu et al., 2022; Luo et al., 2022; Xu et al., 2022), most studies assume that the target is known or manually identified and given as input, a rare scenario in real-world applications where the target is often uncommon, unknown or not explicitly conveyed.
+Also, target annotation is an expensive task in SD (Küçük and Can, 2020).
 In this paper, we focus on a different yet challenging task we refer to as Open-Target Stance Detection (OTSD).
@@
 The target may be mentioned explicitly or not and the stance is predicted with respect to the produced target.
+It is worth noting that in cross-target stance detection (Zhang et al., 2020), while the target is also unseen during training, it typically belongs to domains similar to those in the training set.
+In contrast, in the OTSD task, we mostly consider scenarios where targets from unrelated or different domains may emerge.
+Both tasks fall under the category of ZSSD.
 There has been one notable attempt to partially address this real-world setting of SD.
@@
 1).
-Our empirical study aims to address the following research questions: [RQ#1] How do proprietary and open LLMs—specifically GPT-4o (Achiam et al., 2023), GPT-3.5 (Brown et al., 2020), Llama 3 (Touvron et al., 2023), and Mistral (Jiang et al., 2023)—perform in open TG compared to the TSE when the real target is explicitly or non-explicitly mentioned in the text for stance classification, and how do they compare to each other?
+Our empirical study aims to address the following research questions: [RQ#1] How do proprietary and open LLMs—particularly models from GPT, Gemini, Llama, and Mistral families, —perform in open TG compared to TSE when the real target is explicitly or non-explicitly mentioned in the text, and how do they compare to each other?
 [RQ#2] How do the same LLMs perform compared to TSE in SD (in the context of OTSD) for both explicit and non-explicit cases, and how do they compare to each other?
@@
 3.1 Task Definition
-Given an input text $x$, previous work (Eq.
+Given an input text $x$, previous work Li et al.
+(2023a) (Eq.
 1) generates a target $t^{\prime}$ and maps it to a predefined list of targets $k$, resulting in a mapped target $t$.
 Finally, they detect the stance $y$ given $x$ and $t$.
-OTSD task (Eq.
-2) is to detect the stance $y$ from the input text $x$ and the generated target $t^{\prime}$, eliminating the need for a predefined list $k$.
+OTSD objective (Eq.
+2) is to detect the stance $y$ from the input text $x$ and the generated target $t^{\prime}$, eliminating the need for a predefined list.
 […]
@@
 To address the research questions outlined in §1, our approach to OTSD leverages the zero-shot learning capabilities of LLMs.
-We adopt the "Task Definition" prompting strategy, as it has been shown by Cruickshank and Ng (2023) to outperform other strategies—such as "Task-only", "Context Analyze", "Context Question", "Zero-shot CoT", and "CoDA"—particularly for smaller models (7B-8B parameters) across most datasets.
-Since this is an open-target task, we use a straightforward prompting approach to ensure fairness and generalizability, avoiding tailored prompts for specific targets, or in-context examples, aligning with the zero-shot nature of this task.
+We adopt the "Task Definition" prompting strategy, as demonstrated by Cruickshank and Ng (2023), which either outperforms or performs on par with alternative strategies such as "Chain-of-Thought" (CoT) and "few-shot" across most stance detection datasets.
+To further validate this, we conduct a preliminary experiment with the CoT strategy (detailed in Appendix K), which reveals that CoT underperforms in TG, and performs comparably to "Task Definition" in SD, reinforcing the early findings.
+Additionally, using the "Task Definition" strategy without relying on tailored prompts for specific targets or in-context examples ensures the generalizability of OTSD, aligning with the zero-shot nature of this task.
 We employ two different prompting approaches for TG and SD.
 The prompts and their design justifications can be found in Appendix A.
-TG+SD (Two-Step Approach) In this approach, we first generate the target and then sequentially detect the stance of the text toward the generated target.
-By directing the LLM to focus on one task at a time, this method may reduce the cognitive load on the model.
-TG&SD (Single-Step Joint Approach) This approach combines TG and SD into a single step using a unified prompt.
-By jointly generating the target and stance, the model can develop a more holistic understanding of the relationship between the text, target, and stance.
+TG+SD (Two-Step Approach) In this approach, we first generate the target and then sequentially detect the stance of the text towards the generated target, which directs LLMs to focus on one task at a time.
+TG&SD (Single-Step Joint Approach) This approach unifies TG and SD in a single step using a prompt, enabling the model to better understand the relationship between the text, target, and stance.
 4 Experiments
@@
 Since OTSD is a generative task, we primarily compare our results with TSE-BestGen, despite its reliance on a predefined list of golden targets.
-Additionally, we consider TSE-mapped as the current best score, regardless of task type.
 Details about our hardware settings can be found in Appendix D.
@@
 To ensure a fair comparison between TSE and LLM models, we use the same dataset as TSE’s zero-shot setup.
-VAST (Allaway and McKeown, 2020) is used to compare LLMs across different settings.
-Both TSE and VAST datasets contain three stance classes (Favor, Against, None), and contain 3,000 and 5,100 samples, respectively, in our study.
-VAST, originally a multi-target dataset, is converted to a single-target format (conversion process in Appendix E) to match the scope of our study.
-TSE has 500 samples for each of 6 targets (see Appendix F), while VAST includes 2,145 unique targets.
-We classify samples as explicit or non-explicit by removing stop words and special characters, lemmatizing all words, and checking for target words in the text.
-This results in 1,804 explicit and 1,196 non-explicit samples from TSE, and 3,120 explicit and 1,980 non-explicit samples from VAST (see Appendix F).
-Note that, due to the lack of code and details from TSE for reproducing target generation and selection, we cannot use other SD datasets for benchmarking.
-In our focused contribution, we use both proprietary and non-proprietary LLMs, including GPT-3.5, GPT-4o, Llama-3, and Mistral to address our two research questions (§1).
+In addition to TSE dataset, we include VAST (Allaway and McKeown, 2020) and EZSTANCE (Zhao and Caragea, 2024) to evaluate LLMs across different settings.
+All three datasets contain three stance classes (Favor, Against, None) and include 3000, 5100, and 9313 samples, respectively, in our study.
+To align with the scope of our study, the multi-target VAST and EZSTANCE datasets are converted into single-target formats (details in Appendix E).
+TSE, VAST, and EZSTANCE feature 6, 2145, and 6873 unique targets, respectively.
+These datasets—particularly VAST and EZSTANCE—cover a wide range of domains and unique targets that are rarely used in stance detection, highlighting the practical challenges of the OTSD task.
+Samples are classified as explicit or non-explicit through preprocessing steps, including stop-word and special character removal, word lemmatization, and checking for target words in the text.
+This process results in 1804 explicit and 1196 non-explicit samples from TSE dataset, 3120 explicit and 1980 non-explicit samples from VAST (see Appendix F), and 9313 explicit and 149 non-explicit samples from EZSTANCE.
+Note that due to the lack of code and details from the TSE method for reproducing target generation and selection, we could not use other SD datasets for benchmarking.
+For models, we use both proprietary and non-proprietary LLMs, including GPT-3.5, GPT-4o, Gemini-flash-8B, Gemini-pro, Llama-3-8B, Llama-3-70B, Mistral-small, and Mistral-large (description in Appendix  B).
 4.2 Evaluation Methods
@@
 We find that the BERTweet F1 score consistently and significantly improved with higher-quality targets, in both explicit and non-explicit cases (see Appendix C).
-More importantly, BTSD score shows strong tau ($\tau$) correlation (KENDALL, 1938) of $0.85$ and $0.79$ for TSE and VAST, respectively, with human judgments gathered during experimentation (see the section below).
-Therefore, the BTSD serves as a reliable proxy for evaluating target generation quality.
-Human Evaluation: To evaluate the relevance of the generated targets and further validate the BTSD metric, we conduct a small-scale human evaluation using 500 randomly selected samples (300 explicit and 200 non-explicit) from both datasets, maintaining an equal ratio of the stance labels.
-Three annotators were asked to assess the relevance of the generated targets to the golden ones, classifying them as either 0 (Not Related), 0.5 (Partially Related), or 1 (Completely Related), following the guidelines provided in Appendix K.
-To measure inter-annotator agreement among the three raters, we compute Krippendorff’s $\alpha$ (Krippendorff, 2011), and Fleiss’s $\kappa$ (Fleiss, 1971) for each model, as shown in Appendix Table 6.
-We observe a solid overall agreement of $\alpha$=0.76 and $\kappa$=0.664.
+More importantly, BTSD score shows strong tau ($\tau$) correlation (Kendall, 1938) of $0.85$, $0.74$, and $0.79$ for TSE, VAST, and EZSTANCE respectively, with human judgments gathered during experimentation.
+Therefore, BTSD serves as a reliable proxy for evaluating target quality.
+Human Evaluation: To evaluate the relevance of the generated targets by some models and further validate the BTSD metric, we conducted a small-scale human evaluation on some of the models using 500 randomly selected samples from each dataset, maintaining an equal ratio of the stance labels.
+Three annotators were asked to assess the relevance of the generated targets to the gold ones, classifying them as either 0 (Not Related), 0.5 (Partially Related), or 1 (Completely Related) (details in Appendix J).
+We observed a solid overall annotator agreement of $\alpha$=0.76 (Krippendorff, 2011) and $\kappa$=0.664 (Fleiss, 1971)(Appendix Table 5).
 SemSim: In OTSD, since we do not rely on a predefined list of targets, the generated targets may not precisely match the gold targets but are semantically related.
 Therefore, we measure the semantic similarity between the generated and gold targets using BERT embeddings.
-Detailed steps can be found in Appendix H with human judgement correlation of $\tau$=$0.57$ for TSE, and $\tau$=$0.59$ for VAST.
+Detailed steps can be found in Appendix H with human judgement correlation of $\tau$=$0.57$ for TSE, and $\tau$=$0.59$ for VAST, way below BTSD.
 We use macro-average F1 for the task of SD.
-TG+SD TG&SD TG+SD TG&SD
-Model SS BTSD HE SC SS BTSD HE SC SS BTSD HE SC SS BTSD HE SC
-Explicit Explicit
+TG+SD TG&SD TG+SD TG&SD TG+SD TG&SD
+Model SS BTSD HE SC SS BTSD HE SC SS BTSD HE SC SS BTSD HE SC SS BTSD HE SC SS BTSD HE SC
+Explicit Explicit Explicit
 […]
-TSE-B 0.86 35.8 0.338 37.81 - - - - - - - - - - - -
+TSE-B 0.86 35.8 0.338 37.81 - - - - - - - - - - - - - - - - - - - -
 […]
@@
 […]
-Non-explicit Non-explicit
+Llama-3-8B 0.88 38.31 0.54 43.13 0.87 37.75 0.558 42.1 0.86 41.13 0.467 47.20 0.88 42.45 0.458 46.72 0.84 47.07 0.435 40.51 0.87 48.79 0.441 43.42
 […]
-TSE-B 0.82 29.32 0.215 31 - - - - - - - - - - - -
+[…]
+[…]
+Non-explicit Non-explicit Non-explicit
+[…]
+TSE-B 0.82 29.32 0.215 31 - - - - - - - - - - - - - - - - - - - -
+[…]
+[…]
+[…]
+[…]
 […]
@@
 Although TSE-mapped achieves higher HE score in explicit case, its BTSD score is lower.
-This is because TSE-mapped produces either a perfect match (score=1) or a completely unrelated target (score=0) favouring higher HE score (see Appendix M).
-In contrast, LLMs mostly generate targets that are highly or partially related to golden targets, resulting in fewer irrelevant outputs and higher BTSD, despite slightly lower average HE scores.
+This is because TSE-mapped produces either a perfect match (score=1) or a completely unrelated target (score=0) favouring higher HE score (see Appendix L).
+In contrast, LLMs mostly generate targets that are highly or partially related to gold targets, resulting in fewer irrelevant outputs and higher BTSD, despite slightly lower average HE scores.
 When the target is not explicitly mentioned, LLMs produce higher-quality targets than both TSE models.
-However, the quality drops for LLMs (TSE as well) compared to explicit cases (as per Table 1), with the best-performing LLM getting the HE score slightly lower than $0.5$ (refers to partially related target) on average.
-A manual inspection of GPT-3.5’s 30 incorrect predictions suggests that the lower score in non-explicit cases is due to insufficient context, often resulting from how the TSE test set is constructed.
-The lack of implicit hints or surrounding context leads the model to generate targets based solely on the text provided, causing inaccuracies (see example in Appendix J.1).
+However, the quality drops for LLMs (TSE as well) compared to explicit cases (as per Table 1), with HE scores slightly lower than $0.5$ (refers to partially related target) on average.
+A manual inspection of the lower quality targets suggest that the lower score in non-explicit cases is due to insufficient context of the text.
+The lack of implicit hints or surrounding context leads the model to generate targets based solely on the text provided, causing inaccuracies (example from GPT-3.5 in Appendix I.1).
 When comparing tested proprietary and open LLMs using BTSD and HE, there is no clear advantage in TG (i.e., not all proprietary models outperform open ones, or vice versa).
-We note that GPT-4o consistently produces targets of equal or higher quality than other tested models across the settings, including explicit and non-explicit target mentions, TG+SD, and TG&SD approaches, and across the datasets.
-Among the tested open LLMs, Mistral generates higher quality targets than Llama-3 in most settings.
-Comparing between the approaches, most models perform better in TG&SD across both datasets, for explicit and non-explicit cases.
+Overall, GPT-4o produces targets of equal or higher quality than other tested proprietary models across the approaches, including explicit and non-explicit target mentions.
+Among the tested open LLMs, Mistral-large and Llama-3-70B outperform each other under different settings.
+Comparing between approaches, most models perform better in TG&SD across both datasets, for explicit and non-explicit cases.
 5.2 Stance Detection (SD)
-As shown in Table 1, according to the SC (stance classification), LLMs outperform both TSE models in detecting stances for explicit cases, while performing comparably with TSE models in non-explicit cases (except GPT-4o).
+As shown in Table 1, according to the SC (stance classification), LLMs outperform both TSE models in detecting stances in both explicit and non-explicit cases.
 In general, LLMs perform better in SD when the target is explicitly mentioned in the text compared to when it is not.
 Our manual analysis shows that in non-explicit cases, the input text lacks sufficient context about the target, leading to lower SC scores.
-Overall, the experimented proprietary models (GPT-3.5 and GPT-4o) outperform the open ones in TG&SD in both the explicit and non-explicit cases across the datasets.
-In some cases of TG+SD setting, though, the open LLMs achieve comparatively higher scores.
-Comparing approaches, LLMs tend to perform better when prompted with the holistic TG&SD, though their TG+SD results remain competitive, specifically in non-explicit cases across both datasets.
-Although the quality of the generated targets correlates positively with the SC score in explicit cases (TSE $\tau$=$0.71$, VAST $\tau$=$0.29$), this correlation is negative in non-explicit cases (TSE $\tau$=$-0.14$, VAST $\tau$=$-0.29$).
-We manually investigate the negative correlation (lower SC with higher quality target) in non-explicit cases.
-Indeed, LLMs such GPT-4o may generate highly relevant targets that are closely related to the golden targets but nearly ‘antithetical’ or opposite, leading to a stance reversal in comparison to the golden stance.
-For example, when the golden target is "permits to carry guns", GPT-4o generates "gun control in universities", while GPT-3.5 generates "guns".
-Although GPT-4o’s target is more closely related to the golden target, it represents the opposite stance.
-On the other hand, GPT-3.5’s target, though broader, less accurate and less related, aligns with the stance of golden target.
-This highlights the need for a coherence measure in OTSD that accounts for the target-stance alignment.
+The tested proprietary and open LLMs exhibit competitive performance against each other in both explicit and non-explicit cases for both approaches, with neither showing a clear advantage.
+When comparing approaches, LLMs generally perform better with the holistic TG&SD strategy in both explicit and non-explicit cases, with the exception of Llama-3-8B, and Mistral-small (for non-explicit).
+The quality of the generated target (BTSD) shows a stronger correlation with the SC score in explicit cases (avg.
+$\tau=0.4$) across datasets compared to non-explicit ones (avg.
+$\tau=0.25$), particularly for VAST and EZSTANCE.
+Upon manual investigation, we observed that LLMs such GPT-4o may generate highly relevant targets that are closely related to the gold targets but nearly ‘antithetical’ or opposite, leading to a stance reversal in comparison to the gold stance.
+For example, in VAST, when the gold target is "permit to carry gun", GPT-4o generates "gun control", while GPT-3.5 generates "guns".
+Although GPT-4o’s target is more closely related to the gold target, it conveys the opposite.
+GPT-3.5’s target, though broader, aligns with the stance of gold target.
 6 Conclusion
-We introduce the Open Target Stance Detection (OTSD) task, where targets are neither seen during training nor provided as input, making it more realistic and challenging.
-We examine proprietary and open LLMs in our OTSD task, and propose a metric to evaluate the quality of Target Generation (TG).
-We compare to TSE, a method leveraging a predefined list of targets.
-LLMs outperform it in TG.
-However, LLMs struggle in generating quality targets when these are not explicitly mentioned.
-LLMs also surpass TSE in Stance Detection (SD) when the target is explicitly mentioned while performing comparably in non-explicit cases.
-Finally, a single-step joint prompt approach (TG&SD) of LLMs proves more effective than a two-step approach (TG+SD) for both TG and SD.
+We introduce the Open Target Stance Detection (OTSD) task, where targets are unseen during training and not provided as input, making it both realistic and challenging.
+We evaluate proprietary and open LLMs on OTSD and propose a metric for assessing Target Generation (TG) quality.
+Compared to TSE, which relies on predefined target lists, LLMs perform well in TG but struggle when targets are not explicitly mentioned.
+LLMs also surpass TSE in Stance Detection (SD) for both explicit and non-explicit targets.
+Finally, a single-step joint prompt (TG&SD) outperforms the two-step approach (TG+SD) for both TG and SD.
 Limitation
 The limitations of our work can be listed as follows:
-  • We did not provide a specific coherence measure to evaluate the Open-Target Stance Detection task, which is worth to be investigated in the future work, as the target and the stance toward it are inherently interconnected.
-  • We conducted a thorough investigation of simple prompting for LLMs, inspired from the stance detection literature (§3.2), which has been shown to outperform other strategies, and provides the answers of our research questions already.
-However, there might be still room for improvement through a more detailed examination of different prompting techniques or task-specific pre-training.
-While we present some preliminary results on the chain-of-thought approach in Appendix L, further investigation is needed to fully understand its potential.
-  • Regarding potential data leakage during the pre-training of LLMs, it is possible that these models were exposed to some of our test data, leading to prior knowledge of certain entities.
-However, since the LLMs were not explicitly pre-trained on the same task, we continue to refer to our approach as zero-shot stance detection.
+  • This work has highlighted specific challenges of the proposed OTSD task and measured both target quality and stance detection performance.
+However, our findings underscore the need to improve the measure of stance detection accuracy in OTSD, enabling it to account for the alignment/coherence between the generated target and the stance toward it, given the text and the reference (gold) target.
+  • A potential limitation of our work lies in the possibility of data leakage during the pre-training of LLMs, where the models may have been exposed to some of the test data, thereby acquiring prior knowledge of certain entities.
+Nevertheless, as the LLMs were not explicitly pre-trained on the same task, we maintain that our approach qualifies as zero-shot stance detection.
+  • In this work, we limit our experiments to the OTSD task in English. Future research can explore the OTSD task in a cross-lingual setting.
   • Although we did not claim that the results or findings are true for all LLMs, experimenting with other proprietary LLMs was financially constraining for us.
@@
 To determine the optimal length, we test various target length instructions across all LLMs (§4.1) on 100 randomly selected samples from our dataset (§4.1), aiming to achieve an average length and standard deviation that closely match those of the golden targets.
-Based on our empirical study, we set the maximum target length to 5 for GPT-3.5 and 4 for Llama and Mistral to obtain an average length and standard deviations close to those of gold targets (explicit: $3.78\pm 1.5$; non-explicit 2.66±1.76)2.66$\pm$ 1.76)2.66 ± 1.76 ).
+Based on our empirical study, we set the maximum target length to 5 for GPT and Gemini models and 4 for Llama and Mistral models to obtain an average length and standard deviations close to those of gold targets (explicit: $3.78\pm 1.5$; non-explicit 2.66±1.76)2.66$\pm$ 1.76)2.66 ± 1.76 ).
 TG+SD (Two-Step Approach)
@@
 A target should be the topic on which the tweet is talking.
-The target can be a single word or a phrase, but its maximum length MUST be 4 words.
+The target can be a single word or a phrase, but its maximum length MUST be #n words.
 The output should only be the target, no other words.
@@
 Only return the stance as a single word, and no other text.
-Prompt Justification:
-  • Clarity of Task: The task description explicitly states what the LLM needs to do—generate a target for a given tweet in TG prompt and for SD, it provides a clear and concise definition of stance classification.
-This ensures that the LLM understands the primary objective without ambiguity.
-  • Focus on Relevance: By instructing the model that the target should be the topic of the tweet, it emphasizes the need for relevance, ensuring that the generated target accurately represents the tweet’s content.
-  • Conciseness Requirement: Limiting the target to a maximum of 4 words ensures that the output remains concise and focused, which is crucial for the generated target to be compared fairly with the golden target.
-  • Stance Options Provided: By specifying the possible stances (Favor, Against, None), the prompt sets clear categories for the model to choose from.
-This simplifies the models decision process and ensures the outputs are standardized.
-  • Output Format Specification: Clearly specifying that the output should only be the target, with no additional words or explanations, helps to standardize the outputs, making them easier to evaluate.
-  • Mandating Output Generation: By instructing the model not to leave any output blank and always to provide an answer, the prompt ensures robustness in responses, reducing the chance of missing data.
 TG&SD (Single-Step Joint Approach)
@@
 A target should be the topic on which the tweet is talking.
-The target can be a single word or a phrase, but its maximum length MUST be 4 words.
+The target can be a single word or a phrase, but its maximum length MUST be #n words.
 If the stance is in favor of the target, write FAVOR, if it is against the target write AGAINST and if it is ambiguous, write NONE.
@@
 The output format should be: “‘Target: <target>, Stance: <stance>“‘.
-Prompt Justification:
-  • Integrated Task Definition: The prompt combines the definitions of both target generation and stance classification, ensuring the LLM understands the combined objective of generating a target and determining the stance towards it.
-  • Clear and Concise Instructions: By clearly outlining the steps—first generating a target and then determining the stance towards it—the prompt guides the model through a logical sequence, reducing ambiguity and improving task performance.
-  • Conciseness Requirement: Limiting the target to a maximum of 4 words ensures that the output remains concise and focused, which is crucial for the generated target to be compared fairly with the golden target.
-  • Target Generation Constraints: Limiting the target to a maximum of 4 words ensures that the output remains concise and focused, which is crucial for the generated target to be compared fairly with the golden target.
-  • Stance Classification Options: Providing explicit stance options (Favor, Against, None) simplifies the decision process for the model and ensures consistency in the outputs.
-  • Mandatory Output Generation: By instructing the model to always provide an output and not leave any fields blank, the prompt ensures robustness and completeness in the responses, reducing the chance of missing data.
-  • Output Format Specification: Specifying the output format as ’Target: <target>, Stance: <stance>’ standardizes the responses, making them easier to parse and evaluate.
-  • Focus on Simplicity and Relevance: The prompt emphasizes that only the stance should be returned as a single word and no additional text, which maintains simplicity and relevance in the outputs.
 Appendix B Model Descriptions
@@
 It builds upon the foundations of GPT-4 (Achiam et al., 2023), incorporating improvements in training and fine-tuning for better performance across diverse language tasks.
-Llama-3-8B Instruct
+Gemini-1.5-Flash-8B
+is a decoder-only model with 8 billion parameters, designed for efficiency and speed.
+Part of the Gemini series, it incorporates optimizations tailored for rapid inference without compromising performance on diverse tasks.
+Gemini-1.5-Pro
+is an advanced decoder-only model in the Gemini series, featuring enhanced architectural modifications and fine-tuned parameters to handle complex multi-domain tasks.
+It is instruction-tuned and optimized for professional-grade applications.
+Llama-3-8B-Instruct
 is a decoder-only model in the LLaMA (Touvron et al., 2023) series with 8 billion parameters that follows the architectural principles of its predecessor, Llama-2 (Touvron et al., 2023), but has been instruction-tuned to improve its performance on a variety of tasks.
-Mistral 7B Instruct
+Llama-3-70B-Instruct
+is a decoder-only model in the LLaMA (Touvron et al., 2023) series with 70 billion parameters.
+Building on the architectural advancements of Llama-2 (Touvron et al., 2023), it has been instruction-tuned to excel across a wide range of natural language processing tasks.
+Mistral-small
 (Jiang et al., 2023) is a decoder-only model like Llama-2 (Touvron et al., 2023), but with different architectural features like Grouped-Query Attention and Sliding Window Attention.
 The Mistral model is 7 billion parameters and has been instruction tuned.
+Mistral-Large
+is a decoder-only model in the Mistral family, incorporating advanced features like Grouped-Query Attention and Sliding Window Attention.
+With 13 billion parameters, it builds upon the success of Mistral-small (Jiang et al., 2023) and has been instruction-tuned for improved task generalization.
 Appendix C BTSD Metric Justification
@@
 Table 2 contains the results of the stratified experiments with different measure stance detection F1 score when giving as input the text and a target, varying the quality of this latter.
-The numbers are significantly different.
+All the numbers in Table 2 are significantly different with p-values < 0.05 in both cases.
 Appendix D Hardware Setting
-All of the experiments were run on a computer with Ubuntu 22.04 Linux with x64 CPU with 4 cores, 32 GB of RAM and one NVIDIA V100 GPU.
-Appendix E VAST Dataset Processing
+All of the experiments were run on a computer with Ubuntu 22.04 Linux with x64 CPU with 8 cores, 32 GB of RAM and two NVIDIA A100 GPUs.
+Appendix E Dataset Pre-processing
+VAST
 Originally, the VAST dataset is a multi-target dataset, with multiple targets assigned to each sample text.
@@
 After this conversion, we obtained a total of 5,100 samples (as shown in Appendix Table 4), covering both explicit and non-explicit samples (the filtering process for explicit and non-explicit samples is described in Section 4.1) from the original 18,545 samples.
+EZSTANCE
+The dataset originally contains 47,316 Twitter samples across three stance categories (favor, against, and neutral).
+Since EZSTANCE features multiple targets for some texts, we adapted it into a single-target dataset (one target per text) to better align with our research by following a series of steps.
+The original dataset is organized into two subtasks (target-based and domain-based), with each subtask containing two target types: claim and noun-phrase.
+First, we combine the samples from both target types and remove duplicates.
+For instances where both target types exist, we retain the noun-phrase version, as it poses a greater challenge in generating precise targets.
+Then, we merge the samples from both subtasks and deduplicate them based on their first occurrence.
+This process results in 9,462 unique samples with 6,873 distinct targets.
+We further divide the dataset into explicit and non-explicit cases, as outlined in Section 4.1, yielding 9,313 explicit and 149 non-explicit samples, where the target is mentioned explicitly or implicitly in the text, respectively.
 Appendix F Dataset Samples
@@
 Note that $t$ is the golden target here.
+During the test phase, initially we considered three random seeds, but finally reported the scores with a single seed, as the score variance was low.
 Appendix H SemSim Details
-Specifically, we employ the bert-base-uncased (Devlin et al., 2018) model to obtain contextual encoded representations of the target words.
+We employ the bert-base-uncased (Devlin et al., 2018) model to obtain contextual encoded representations of the target words.
 After performing a mean pooling operation on these representations, we apply cosine similarity (Singhal, 2001) to the encoded sequences to obtain the similarity score.
-Appendix I OTSD with EZSTANCE
-We explore the OTSD task using EZSTANCE (Zhao and Caragea, 2024), the largest zero-shot stance detection (ZSSD) dataset, to address how proprietary and open LLMs compare in both target generation (TG) and stance detection (SD) tasks when the target is explicitly or implicitly mentioned in the text.
-Given the open-target nature of our task, EZSTANCE was selected for its diverse range of domains and unique target set, reflecting real-world scenarios.
-I.1 Dataset Processing
-The dataset originally contains 47,316 Twitter samples across three stance categories (favor, against, and neutral).
-Since EZSTANCE features multiple targets for some texts, we adapted it into a single-target dataset (one target per text) to better align with our research by following a series of steps.
-The original dataset is organized into two subtasks (target-based and domain-based), with each subtask containing two target types: claim and noun-phrase.
-First, we combine the samples from both target types and remove duplicates.
-For instances where both target types exist, we retain the noun-phrase version, as it poses a greater challenge in generating precise targets.
-Then, we merge the samples from both subtasks and deduplicate them based on their first occurrence.
-This process results in 9,462 unique samples with 6,873 distinct targets.
-We further divide the dataset into explicit and non-explicit cases, as outlined in Section 4.1, yielding 9,313 explicit and 149 non-explicit samples, where the target is mentioned explicitly or implicitly in the text, respectively.
-I.2 Result Analysis
-As shown in Table 5, LLMs face challenges in target generation for non-explicit samples compared to when the target is explicitly mentioned, based on BTSD and HE scores.
-When comparing proprietary and open models, consistent with our earlier findings (§5.1), no single LLM consistently outperforms the others.
-While GPT-4o tends to generate higher-quality targets in the TG+SD approach, Mistral surpasses other models in some TG&SD cases.
-Across both explicit and non-explicit samples, all models perform noticeably better using the TG&SD approach.
-For stance detection, as shown in 5, similar to our findings with TSE and VAST in Section 5.2, LLMs perform better in SC when the target is explicitly mentioned.
-When comparing proprietary and open models, GPT-4o and Mistral both demonstrate competitive performance across approaches.
-Overall, LLMs consistently achieve better stance detection in the TG&SD approach.
-TG+SD TG&SD
-Model SS BTSD HE SC SS BTSD HE SC
-Explicit
-[…]
-[…]
-[…]
-[…]
-Non-explicit
-GPT-3.5 0.76 38.78 0.377 36.69 0.81 41.89 0.402 38.97
-[…]
-[…]
-[…]
-Table 5: Comparison of TG and SD performance on EZSTANCE dataset (Zhao and Caragea, 2024) among the LLMs using various metrics: TG is evaluated with SS (SemSim), BTSD (%), and HE (Human Evaluation) scores, while SC (%) measures SD performance.
-Metric-wise best results within a specific setting are underlined, and better results of a model across TG+SD and TG&SD for each dataset are in bold.
-Appendix J Explanation with Samples in TG and SD
-J.1 Target Generation
+Appendix I Explanation with Samples in TG and SD
+I.1 Target Generation
 For instance, in the post "I mean, this reads to me like anyone who’s an alcoholic doesn’t go to heaven…", GPT-3.5 generated "Sinful behavior and heaven’s eligibility" as the target instead of the gold target "gay rights".
-J.2 Stance Detection
+I.2 Stance Detection
 For instance, in the explicit sample text, "Q.
@@
 This mistake reflects the model’s need for broader context to establish the correct link.
-Appendix K Human Annotator Agreement
+Appendix J Human Annotation Process
+Three graduate students, specializing in NLP, participated voluntarily as annotators in the evaluation of generated targets.
+They were chosen for their expertise and familiarity with the stance detection task, enabling them to effectively assess the relevance of the generated targets compared to the golden ones.
+First, we provided them with an annotation guideline (detailed in §J.1) that included descriptions of the datasets, an overview of the task, and a descriptive scoring scale accompanied by examples.
+Each annotator then independently scored the targets generated by different LLMs for 500 randomly selected samples (300 explicit and 200 non-explicit).
+To assess inter-annotator agreement, we calculated Krippendorff’s $\alpha$ (Krippendorff, 2011) and Fleiss’s $\kappa$ (Fleiss, 1971), with the results presented in Table 5.
+The final score for each generated target was determined based on the majority vote among the annotators.
+In cases of a tie, the median score was used.
+J.1 Annotation Guideline
 TSE VAST
@@
 The reliability score is 1 for TSE-M as it classifies between the golden targets instead of generating.
-Dataset Description: The file includes a Tweet column with various Twitter posts.
+Dataset Description: The file includes a Text column with various Twitter posts or news comments.
 The Gold Target column represents the true target, indicating the primary topic or issue the tweet addresses, which in turn informs the stance noted in the Gold Stance column.
-For each tweet, there are eight generated targets (in columns T1 through T8) produced by different models, whose names remain anonymous for unbiased annotation.
+For each tweet, there are twenty-four generated targets (in columns T1 through T24) produced by different models, whose names remain anonymous for unbiased annotation.
 Additionally, there are empty cells (e.g., in the T1 Score column) designated for storing human evaluation scores corresponding to each generated target.
-Definition of a Relevant Target: A generated target in the context of the target generation task is considered relevant if it accurately captures or closely aligns with the main topic, issue, or entity identified by the golden target in a given tweet.
-It should reflect the key subject matter or concern that the golden target represents, maintaining coherence with the overall context and meaning of the tweet.
-Task Description: Your task is to assess the quality of each generated target by evaluating its relevance to the golden target in the tweet.
+Definition of a Relevant Target: A generated target in the context of the target generation task is considered relevant if it accurately captures or closely aligns with the main topic, issue, or entity identified by the golden target in a given text.
+It should reflect the key subject matter or concern that the golden target represents, maintaining coherence with the overall context and meaning of the text.
+Task Description: Your task is to assess the quality of each generated target by evaluating its relevance to the golden target in the text.
 You will do this by answering the following question and assigning a score to each generated target.
 Scoring criteria and examples are provided below.
-“How closely does the generated target relate to the golden target in the tweet?”
+“How closely does the generated target relate to the golden target in the text?”
 Scoring Scale: 0 to 1 (Low to High)
@@
     T2: superstition, T2 Score: 0.
-Metrics TG+SD TG&SD
-Ex N-Ex Ex N-Ex
-BTSD-TSE 38.47 30.73 37.88 32.14
-BTSD-VAST 41.10 41.54 41.34 40.38
-SC-TSE 47.36 38.11 49.14 36.00
-SC-VAST 48.47 46.46 49.00 45.41
-Table 7: Caption goes here TG (evaluated by BTSD) and SD (evaluated by SC) performance using COT prompting on both TSE and VAST datasets in both Explicit and Non-explicit cases across the two approaches: TG+SD, TG&SD.
-Ex and N-Ex represent Explicit and Non-explicit cases, respectively.
-Appendix L COT Initial Result
-We initiate an exploration of alternative prompting techniques by employing the chain-of-thought (COT) prompting strategy for both target generation (TG) and stance detection (SD) tasks, across two approaches: TG+SD and TG&SD.
-As an initial phase, in Table 7, we assess target quality using the BTSD metric and evaluate stance detection performance with the SC metric, applied to both the TSE and VAST datasets for explicit and non-explicit cases.
-Appendix M Human Annotation Score Distribution
+TG+SD TG&SD
+Metrics TSE VAST EZSTANCE TSE VAST EZSTANCE
+Explicit
+BTSD (GPT-3.5) 38.47 (38.43) 41.10 (41.67) 47.60 (49.27) 37.88 (39.60) 41.34 (44.25) 48.98 (49.70)
+BTSD (GPT-4o) 40.63 (41.55) 41.59 (42.69) 48.12 (49.70) 41.31 (41.92) 42.33 (44.25) 48.76 (50.69)
+SC (GPT-3.5) 37.91 (42.68) 47.47 (47.21) 40.12 (39.10) 44.86 (47.61) 47.00 (48.48) 40.74 (40.63)
+SC (GPT-4o) 47.36 (44.78) 41.17 (40.77) 43.79 (45.93) 49.14 (46.83) 47.84 (49.38) 45.68 (46.22)
+Non-explicit
+BTSD (GPT-3.5) 30.73 (33.10) 41.54 (38.10) 40.76 (38.78) 32.14 (31.32) 40.38 (38.55) 40.54 (41.89)
+BTSD (GPT-4o) 32.72 (35.14) 41.71 (38.55) 41.34 (42.78) 33.44 (36.12) 41.92 (39.84) 43.90 (45.70)
+SC (GPT-3.5) 31.88 (32.66) 43.41 (45.54) 37.76 (36.69) 35.46 (33.94) 44.46 (45.80) 38.10 (38.97)
+SC (GPT-4o) 36.00 (36.39) 38.27 (39.92) 38.34 (38.50) 38.11 (37.5) 42.81 (43.84) 37.54 (39.26)
+Table 6: Preliminary results using the CoT prompting strategy with GPT-3.5 and GPT-4o models on three datasets (TSE, VAST, EZSTANCE), evaluated for both Target Generation (TG) using BTSD and Stance Detection (SD) using SC.
+Results in the parenthesis are from Table 1 following the "Task Definition" strategy.
+The results are presented for both Explicit (Ex) and Non-explicit (N-Ex) cases across two approaches: TG+SD and TG&SD.
+Appendix K Chain-of-Thought (CoT) Results
+We conduct a preliminary experiment using the CoT prompting strategy across both approaches: TG+SD and TG&SD in Table 6.
+Evaluating target generation and stance detection, we test the performance of CoT with GPT-3.5 and GPT-4o models and compare it to the scores achieved using the "Task Definition" strategy, as reported in Table 1.
+The results show that, based on the BTSD score, the "Task Definition" strategy generally outperforms CoT in target generation.
+However, for stance detection, evaluated using SC scores, the two strategies alternate in outperforming each other depending on the specific settings.
+The CoT prompt used in this experiment is as follows:
+Prompt (TG):
+Step 1: You will be provided with a text, think step by step, and explain what is the target of the text. A target should be the topic on which the text is talking.
+Step 2: Therefore, based on your explanation, what is the final target of the text?
+The target can be a single word or a phrase, but its maximum length MUST be 5 words.
+The output should only be the target, NO OTHER WORDS or EXPLANATION.
+Prompt (SD):
+Step 1: Stance classification is the task of determining the expressed or implied opinion, or stance, of a statement toward a certain, specified target.
+Analyze the following text, think step by step, and explain the stance (favor, against, or none) of the text towards the provided target.
+Step 2: Therefore, based on your explanation, what is the final stance?
+If the stance is in favor of the target, write FAVOR, if it is against the target write AGAINST and if it is ambiguous, write NONE.
+Only return the stance as a single word, NO OTHER WORDS or EXPLANATION.
+Prompt (TG, SD):
+Step 1: Stance classification is the task of determining the expressed or implied opinion, or stance, of a statement toward a certain, specified target.
+Analyze the following text, think step by step, and explain what is the target of the text.
+A target should be the topic on which the text is talking.
+Based on the target explanation, think step by step, and explain the stance (favor, against, or none) of the text toward the target.
+Step 2: Therefore, based on your explanation of the target and the stance, generate the target of the given text and determine the stance toward the generated target.
+The target can be a single word or a phrase, but its maximum length MUST be 5 words.
+If the stance is in favor of the target, write FAVOR; if it is against the target, write AGAINST; if it is ambiguous, write NONE.
+Only return the stance as a single word, NO OTHER WORDS or EXPLANATION.
+The final output format MUST be: “‘Target: <target>, Stance: <stance>’ ”.
+Appendix L Human Annotation Score Distribution
 After conducting the human annotation for target quality assessment, we plot the score distribution provided by the annotators, as shown in Figure 1.
@@
 In cases of a tie, the average score of 0.5 is assigned.
-The details on human annotator guideline are provided in Appendix K, and the annotator agreement is in Appendix Table 6.
-Appendix N Vsiual Comparison of TG and SD Performance
-To better illustrate the data in Table 1, we have plotted Figures 2 and 3, which represent the TG and SD performance, respectively.
-These figures provide a clearer visualization of the conclusions discussed in Sections 5.1 and 5.2.
-Figure 2: Comparison of TG performance of TSE and LLMs, as well as the performance among LLMs across both the TSE and VAST datasets, using various metrics: SS (Semantic Similarity), BTSD (%), and HE (Human Evaluation).
-The graphs are generated based on the data in Table 1.
-Figure 3: Comparison of SD performance of TSE and LLMs, as well as the performance among LLMs across both the TSE and VAST datasets, using SC (stance classification) score.
-The graphs are generated based on the data in Table 1.
+The details on human annotator guideline are provided in Appendix J.1, and the annotator agreement is in Appendix Table 5.

### v6

@@
 We adopt the "Task Definition" prompting strategy, as demonstrated by Cruickshank and Ng (2023), which either outperforms or performs on par with alternative strategies such as "Chain-of-Thought" (CoT) and "few-shot" across most stance detection datasets.
-To further validate this, we conduct a preliminary experiment with the CoT strategy (detailed in Appendix K), which reveals that CoT underperforms in TG, and performs comparably to "Task Definition" in SD, reinforcing the early findings.
+To further validate this, we conduct a preliminary experiment with the CoT strategy (detailed in Appendix J), which reveals that CoT underperforms in TG, and performs comparably to "Task Definition" in SD, reinforcing the early findings.
 Additionally, using the "Task Definition" strategy without relying on tailored prompts for specific targets or in-context examples ensures the generalizability of OTSD, aligning with the zero-shot nature of this task.
+This approach first names the task and then provides a concise, self-contained definition of what that task entails—its purpose, inputs, and expected outputs—before asking the model to perform it.
 We employ two different prompting approaches for TG and SD.
@@
 Details about our hardware settings can be found in Appendix D.
+Dataset Source #samples #unique -targets #stance
+Explicit Non-explicit
+TSE Tweets 1,804 1,196 6 3
+VAST News comments 3,120 1,980 2,145 3
+EZSTANCE Tweets 9,313 149 6,873 3
+Table 1: Statistical overview of the datasets utilized for the OTSD task.
 4.1 Dataset and Model
@@
 In addition to TSE dataset, we include VAST (Allaway and McKeown, 2020) and EZSTANCE (Zhao and Caragea, 2024) to evaluate LLMs across different settings.
-All three datasets contain three stance classes (Favor, Against, None) and include 3000, 5100, and 9313 samples, respectively, in our study.
+As shown in Table 1, all three datasets contain three stance classes (Favor, Against, None) and include 3000, 5100, and 9313 samples, respectively, in our study.
 To align with the scope of our study, the multi-target VAST and EZSTANCE datasets are converted into single-target formats (details in Appendix E).
@@
 For models, we use both proprietary and non-proprietary LLMs, including GPT-3.5, GPT-4o, Gemini-flash-8B, Gemini-pro, Llama-3-8B, Llama-3-70B, Mistral-small, and Mistral-large (description in Appendix  B).
-4.2 Evaluation Methods
-BTSD: To measure the generated target quality, we fine-tune the BERTweet model (Nguyen et al., 2020) following the same setup as TSE (fine-tuning details in Appendix G).
-The SD score (F1-macro) is used as the evaluator for the generated targets.
-To assess the effectiveness of this metric, we experiment with different levels of target quality by modifying or removing words from the golden targets, substituting them with incorrect targets, and selecting random words from the vocabulary as targets.
-We find that the BERTweet F1 score consistently and significantly improved with higher-quality targets, in both explicit and non-explicit cases (see Appendix C).
-More importantly, BTSD score shows strong tau ($\tau$) correlation (Kendall, 1938) of $0.85$, $0.74$, and $0.79$ for TSE, VAST, and EZSTANCE respectively, with human judgments gathered during experimentation.
-Therefore, BTSD serves as a reliable proxy for evaluating target quality.
-Human Evaluation: To evaluate the relevance of the generated targets by some models and further validate the BTSD metric, we conducted a small-scale human evaluation on some of the models using 500 randomly selected samples from each dataset, maintaining an equal ratio of the stance labels.
-Three annotators were asked to assess the relevance of the generated targets to the gold ones, classifying them as either 0 (Not Related), 0.5 (Partially Related), or 1 (Completely Related) (details in Appendix J).
-We observed a solid overall annotator agreement of $\alpha$=0.76 (Krippendorff, 2011) and $\kappa$=0.664 (Fleiss, 1971)(Appendix Table 5).
-SemSim: In OTSD, since we do not rely on a predefined list of targets, the generated targets may not precisely match the gold targets but are semantically related.
-Therefore, we measure the semantic similarity between the generated and gold targets using BERT embeddings.
-Detailed steps can be found in Appendix H with human judgement correlation of $\tau$=$0.57$ for TSE, and $\tau$=$0.59$ for VAST, way below BTSD.
-We use macro-average F1 for the task of SD.
 TG+SD TG&SD TG+SD TG&SD TG+SD TG&SD
@@
 […]
-Llama-3-8B 0.88 38.31 0.54 43.13 0.87 37.75 0.558 42.1 0.86 41.13 0.467 47.20 0.88 42.45 0.458 46.72 0.84 47.07 0.435 40.51 0.87 48.79 0.441 43.42
+Llama-3-8B 0.88 38.31 0.54 43.13 0.87 37.75 0.558 42.10 0.86 41.13 0.467 47.20 0.88 42.45 0.458 46.72 0.84 47.07 0.435 40.51 0.87 48.79 0.441 43.42
 […]
@@
 […]
-TSE-B 0.82 29.32 0.215 31 - - - - - - - - - - - - - - - - - - - -
+TSE-B 0.82 29.32 0.215 31.00 - - - - - - - - - - - - - - - - - - - -
 […]
@@
 […]
-[…]
+Mistral-small 0.83 30.18 0.386 33.61 0.83 34.12 0.495 30.47 0.80 39.75 0.405 43.34 0.81 40.05 0.447 41.87 0.77 38.71 0.400 41.36 0.78 46.63 0.414 39.41
 […]
@@
 Metric-wise best results within a specific setting are underlined, and better results of a model across TG+SD and TG&SD for each dataset are in bold.
+4.2 Evaluation Methods
+BTSD: To measure the generated target quality, we fine-tune the BERTweet model (Nguyen et al., 2020) following the same setup as TSE.
+We consider 4 datasets combined i.e.
+SemEval (Mohammad et al., 2016b), AM (Stab et al., 2018), COVID-19 (Glandt et al., 2021), P-Stance (Li and Caragea, 2021), containing 19 targets, splitting samples by (70%-train, 10%-dev, 20%-test), maintaining equal distribution from each of the targets.
+To assess the quality of the generated targets, we train a stance classifier model (referred as BTSD) as follows.
+Given a text $x$ and a target $t$, we first formulate the input as a sequence $s=[[CLS]t[SEP]x]$ where $[CLS]$ is a token that encodes the sentence and $[SEP]$ is used to separate the sentence $x$ and the target $t$.
+Then the representation of $[CLS]$ token is used to predict the stance toward target $t$.
+Note that $t$ is the golden target here.
+During the test phase, initially we considered three random seeds, but finally reported the scores with a single seed, as the score variance was low.
+The SD score (F1-macro) is used as the evaluator for the generated targets.
+To assess the effectiveness of this metric, we experiment with different levels of target quality by modifying or removing words from the golden targets, substituting them with incorrect targets, and selecting random words from the vocabulary as targets.
+We find that the BERTweet F1 score consistently and significantly improved with higher-quality targets, in both explicit and non-explicit cases (see Appendix C).
+More importantly, BTSD score shows strong tau ($\tau$) correlation (Kendall, 1938) of $0.85$, $0.74$, and $0.79$ for TSE, VAST, and EZSTANCE respectively, with human judgments gathered during experimentation.
+Therefore, BTSD serves as a reliable proxy for evaluating target quality.
+Human Evaluation: To evaluate the relevance of the generated targets by some models and further validate the BTSD metric, we conducted a small-scale human evaluation on some of the models using 500 randomly selected samples from each dataset, maintaining an equal ratio of the stance labels.
+Three annotators were asked to assess the relevance of the generated targets to the gold ones, classifying them as either 0 (Not Related), 0.5 (Partially Related), or 1 (Completely Related) (details in Appendix I).
+We observed a solid overall annotator agreement of $\alpha$=0.76 (Krippendorff, 2011) and $\kappa$=0.664 (Fleiss, 1971)(Appendix Table 6).
+SemSim: In OTSD, since we do not rely on a predefined list of targets, the generated targets may not precisely match the gold targets but are semantically related.
+Therefore, we measure the semantic similarity between the generated and gold targets using BERT embeddings.
+Detailed steps can be found in Appendix G with human judgement correlation of $\tau$=$0.57$ for TSE, and $\tau$=$0.59$ for VAST, way below BTSD.
+We use macro-average F1 for the task of SD.
 5 Result Analysis
@@
 Although TSE-mapped achieves higher HE score in explicit case, its BTSD score is lower.
-This is because TSE-mapped produces either a perfect match (score=1) or a completely unrelated target (score=0) favouring higher HE score (see Appendix L).
+This is because TSE-mapped produces either a perfect match (score=1) or a completely unrelated target (score=0) favouring higher HE score (see Appendix K).
 In contrast, LLMs mostly generate targets that are highly or partially related to gold targets, resulting in fewer irrelevant outputs and higher BTSD, despite slightly lower average HE scores.
@@
 A manual inspection of the lower quality targets suggest that the lower score in non-explicit cases is due to insufficient context of the text.
-The lack of implicit hints or surrounding context leads the model to generate targets based solely on the text provided, causing inaccuracies (example from GPT-3.5 in Appendix I.1).
+The lack of implicit hints or surrounding context leads the model to generate targets based solely on the text provided, causing inaccuracies (example from GPT-3.5 in Appendix H.1).
 When comparing tested proprietary and open LLMs using BTSD and HE, there is no clear advantage in TG (i.e., not all proprietary models outperform open ones, or vice versa).
@@
 Following TSE, our targets are Creationism (Somasundaran and Wiebe, 2010), Gay Rights (Somasundaran and Wiebe, 2010), Climate Change is a Concern (Mohammad et al., 2016a), MeToo Movement (Gautam et al., 2020), Merger of Disney and Fox (Conforti et al., 2020), and Lockdown in New York State (Miao et al., 2020).
-Samples from TSE and VAST in both explicit and non-explicit settings are given in Table 3 and 4, respectively.
+Samples from TSE and VAST in both explicit and non-explicit settings are given in Table 4 and 5, respectively.
 Target Text Stance
@@
 Table 5: Samples from VAST dataset with few different targets in both explicit and non-explicit settings.
-Appendix G BTSD Fine-tuning
-Following TSE, we consider 4 datasets combined i.e.
-SemEval (Mohammad et al., 2016b), AM (Stab et al., 2018), COVID-19 (Glandt et al., 2021), P-Stance (Li and Caragea, 2021), containing 19 targets, splitting samples by (70%-train, 10%-dev, 20%-test), maintaining equal distribution from each of the targets.
-To assess the quality of the generated targets, we train a stance classifier model (referred as BTSD) as follows.
-Given a text $x$ and a target $t$, we first formulate the input as a sequence $s=[[CLS]t[SEP]x]$ where $[CLS]$ is a token that encodes the sentence and $[SEP]$ is used to separate the sentence $x$ and the target $t$.
-Then the representation of $[CLS]$ token is used to predict the stance toward target $t$.
-Note that $t$ is the golden target here.
-During the test phase, initially we considered three random seeds, but finally reported the scores with a single seed, as the score variance was low.
-Appendix H SemSim Details
+Appendix G SemSim Details
 We employ the bert-base-uncased (Devlin et al., 2018) model to obtain contextual encoded representations of the target words.
 After performing a mean pooling operation on these representations, we apply cosine similarity (Singhal, 2001) to the encoded sequences to obtain the similarity score.
-Appendix I Explanation with Samples in TG and SD
-I.1 Target Generation
+Appendix H Explanation with Samples in TG and SD
+H.1 Target Generation
 For instance, in the post "I mean, this reads to me like anyone who’s an alcoholic doesn’t go to heaven…", GPT-3.5 generated "Sinful behavior and heaven’s eligibility" as the target instead of the gold target "gay rights".
-I.2 Stance Detection
+H.2 Stance Detection
 For instance, in the explicit sample text, "Q.
@@
 This mistake reflects the model’s need for broader context to establish the correct link.
-Appendix J Human Annotation Process
+Appendix I Human Annotation Process
 Three graduate students, specializing in NLP, participated voluntarily as annotators in the evaluation of generated targets.
 They were chosen for their expertise and familiarity with the stance detection task, enabling them to effectively assess the relevance of the generated targets compared to the golden ones.
-First, we provided them with an annotation guideline (detailed in §J.1) that included descriptions of the datasets, an overview of the task, and a descriptive scoring scale accompanied by examples.
+First, we provided them with an annotation guideline (detailed in §I.1) that included descriptions of the datasets, an overview of the task, and a descriptive scoring scale accompanied by examples.
 Each annotator then independently scored the targets generated by different LLMs for 500 randomly selected samples (300 explicit and 200 non-explicit).
@@
 In cases of a tie, the median score was used.
-J.1 Annotation Guideline
+I.1 Annotation Guideline
 TSE VAST
@@
 The results are presented for both Explicit (Ex) and Non-explicit (N-Ex) cases across two approaches: TG+SD and TG&SD.
-Appendix K Chain-of-Thought (CoT) Results
+Appendix J Chain-of-Thought (CoT) Results
 We conduct a preliminary experiment using the CoT prompting strategy across both approaches: TG+SD and TG&SD in Table 7.
@@
 The final output format MUST be: “‘Target: <target>, Stance: <stance>’ ”.
-Appendix L Human Annotation Score Distribution
+Appendix K Human Annotation Score Distribution
 After conducting the human annotation for target quality assessment, we plot the score distribution provided by the annotators, as shown in Figure 1.
@@
 In cases of a tie, the average score of 0.5 is assigned.
-The details on human annotator guideline are provided in Appendix J.1, and the annotator agreement is in Appendix Table 5.
+The details on human annotator guideline are provided in Appendix I.1, and the annotator agreement is in Appendix Table 6.

### v7

@@
 However, LLMs struggle in both target generation and stance detection when the target is not explicit.
-[^1] Dataset and code are available here: anonymous.4open.science/r/opentarget-5521.
+[^1] Dataset and code are available here: https://github.com/AbuUbaida/opentarget.
 […]
@@
 We adopt the "Task Definition" prompting strategy, as demonstrated by Cruickshank and Ng (2023), which either outperforms or performs on par with alternative strategies such as "Chain-of-Thought" (CoT) and "few-shot" across most stance detection datasets.
-To further validate this, we conduct a preliminary experiment with the CoT strategy (detailed in Appendix J), which reveals that CoT underperforms in TG, and performs comparably to "Task Definition" in SD, reinforcing the early findings.
+To further validate this, we conduct a preliminary experiment with the CoT strategy (detailed in Appendix I), which reveals that CoT underperforms in TG, and performs comparably to "Task Definition" in SD, reinforcing the early findings.
 Additionally, using the "Task Definition" strategy without relying on tailored prompts for specific targets or in-context examples ensures the generalizability of OTSD, aligning with the zero-shot nature of this task.
@@
 Human Evaluation: To evaluate the relevance of the generated targets by some models and further validate the BTSD metric, we conducted a small-scale human evaluation on some of the models using 500 randomly selected samples from each dataset, maintaining an equal ratio of the stance labels.
-Three annotators were asked to assess the relevance of the generated targets to the gold ones, classifying them as either 0 (Not Related), 0.5 (Partially Related), or 1 (Completely Related) (details in Appendix I).
+Three annotators were asked to assess the relevance of the generated targets to the gold ones, classifying them as either 0 (Not Related), 0.5 (Partially Related), or 1 (Completely Related) (details in Appendix H).
 We observed a solid overall annotator agreement of $\alpha$=0.76 (Krippendorff, 2011) and $\kappa$=0.664 (Fleiss, 1971)(Appendix Table 6).
@@
 A manual inspection of the lower quality targets suggest that the lower score in non-explicit cases is due to insufficient context of the text.
-The lack of implicit hints or surrounding context leads the model to generate targets based solely on the text provided, causing inaccuracies (example from GPT-3.5 in Appendix H.1).
+The lack of implicit hints or surrounding context leads the model to generate targets based solely on the text provided, causing inaccuracies.
+For instance, for the post "I mean, this reads to me like anyone who’s an alcoholic doesn’t go to heaven…", GPT-3.5 generated the target "Sinful behavior and heaven’s eligibility" instead of the correct gold target "gay rights".
 When comparing tested proprietary and open LLMs using BTSD and HE, there is no clear advantage in TG (i.e., not all proprietary models outperform open ones, or vice versa).
@@
 GPT-3.5’s target, though broader, aligns with the stance of gold target.
+Another example, in the explicit sample text, "Q.
+from audience: Malcom X said the most disrespected person in America is the black woman, has #MeToo changed this?
+Gender has been constructed in the image of the white woman, the idea of a level playing field, white women are often in better positions, says @HeidiMirza #CHevents", GPT-3.5 generated the target "Gender roles", which, while related, fails to capture the specific stance-related target "MeToo movement".
+This mistake reflects the model’s need for broader context to establish the correct link.
+A potential way to address this issue would be leveraging an LLM as a judge to assess the stance alignment.
+We explore the feasibility of this idea by conducting a small-scale experiment in Appendix J.
 6 Conclusion
@@
 However, our findings underscore the need to improve the measure of stance detection accuracy in OTSD, enabling it to account for the alignment/coherence between the generated target and the stance toward it, given the text and the reference (gold) target.
+To investigate the feasibility of a potential solution, we carried out a preliminary experiment using an LLM as an automated judge of stance-target coherence.
+The LLM’s assessments showed strong correlation with human judgments, suggesting that LLMs can serve as a reliable component in the development of coherence-based evaluation metrics.
+Full details of this pilot study are provided in Appendix J.
+  • In this work, we focus on identifying a single target from each text, even though a given text may contain multiple targets, whether explicitly stated or implied.
   • A potential limitation of our work lies in the possibility of data leakage during the pre-training of LLMs, where the models may have been exposed to some of the test data, thereby acquiring prior knowledge of certain entities.
@@
   • In this work, we limit our experiments to the OTSD task in English. Future research can explore the OTSD task in a cross-lingual setting.
-  • Although we did not claim that the results or findings are true for all LLMs, experimenting with other proprietary LLMs was financially constraining for us.
+  • Although our findings were evaluated across a diverse set of eight LLMs spanning four different families—covering both small and large models, as well as proprietary and open-source ones—future work could benefit from exploring a broader range of models, particularly Large Reasoning Models (LRMs), which present an especially promising direction.
+Ethics Statement
+The datasets used for fine-tuning and zero-shot experiments are publicly available and anonymized, with no associated data protection concerns.
+Additionally, the human annotators involved in the evaluation were informed about the intended use of their contributions and participated voluntarily without compensation.
+The only data collected consisted of their annotation ratings.
+Acknowledgments
+This research was partially supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) under the Grant No.
+RGPIN-2022-04789, and by the Digital Research Alliance of Canada (https://www.alliancecan.ca/en).
+We would like to thank the annotators and the reviewers for their valuable feedback.
 Appendix A Prompting Details
@@
 Samples from TSE and VAST in both explicit and non-explicit settings are given in Table 4 and 5, respectively.
+The demonstration of samples can also be found in the accompanying GitHub repository [^3] GitHub link for sample demonstration: https://github.com/AbuUbaida/opentarget..
 Target Text Stance
@@
 After performing a mean pooling operation on these representations, we apply cosine similarity (Singhal, 2001) to the encoded sequences to obtain the similarity score.
-Appendix H Explanation with Samples in TG and SD
-H.1 Target Generation
-For instance, in the post "I mean, this reads to me like anyone who’s an alcoholic doesn’t go to heaven…", GPT-3.5 generated "Sinful behavior and heaven’s eligibility" as the target instead of the gold target "gay rights".
-H.2 Stance Detection
-For instance, in the explicit sample text, "Q.
-from audience: Malcom X said the most disrespected person in America is the black woman, has #MeToo changed this?
-Gender has been constructed in the image of the white woman, the idea of a level playing field, white women are often in better positions, says @HeidiMirza #CHevents", GPT-3.5 generated the target "Gender roles", which, while related, fails to capture the specific stance-related target "MeToo movement".
-This mistake reflects the model’s need for broader context to establish the correct link.
-Appendix I Human Annotation Process
+Appendix H Human Annotation Process
 Three graduate students, specializing in NLP, participated voluntarily as annotators in the evaluation of generated targets.
 They were chosen for their expertise and familiarity with the stance detection task, enabling them to effectively assess the relevance of the generated targets compared to the golden ones.
-First, we provided them with an annotation guideline (detailed in §I.1) that included descriptions of the datasets, an overview of the task, and a descriptive scoring scale accompanied by examples.
+First, we provided them with an annotation guideline (detailed in §H.1) that included descriptions of the datasets, an overview of the task, and a descriptive scoring scale accompanied by examples.
 Each annotator then independently scored the targets generated by different LLMs for 500 randomly selected samples (300 explicit and 200 non-explicit).
@@
 In cases of a tie, the median score was used.
-I.1 Annotation Guideline
+H.1 Annotation Guideline
 TSE VAST
@@
 Explicit
-BTSD (GPT-3.5) 38.47 (38.43) 41.10 (41.67) 47.60 (49.27) 37.88 (39.60) 41.34 (44.25) 48.98 (49.70)
-BTSD (GPT-4o) 40.63 (41.55) 41.59 (42.69) 48.12 (49.70) 41.31 (41.92) 42.33 (44.25) 48.76 (50.69)
-SC (GPT-3.5) 37.91 (42.68) 47.47 (47.21) 40.12 (39.10) 44.86 (47.61) 47.00 (48.48) 40.74 (40.63)
-SC (GPT-4o) 47.36 (44.78) 41.17 (40.77) 43.79 (45.93) 49.14 (46.83) 47.84 (49.38) 45.68 (46.22)
+BTSD (GPT-3.5) 38.47 (38.43$\uparrow$ 41.10 (41.67$\downarrow$ 47.60 (49.27$\downarrow$ 37.88 (39.60$\downarrow$ 41.34 (44.25$\downarrow$ 48.98 (49.70$\downarrow$
+BTSD (GPT-4o) 40.63 (41.55$\downarrow$ 41.59 (42.69$\downarrow$ 48.12 (49.70$\downarrow$ 41.31 (41.92$\downarrow$ 42.33 (44.25$\downarrow$ 48.76 (50.69$\downarrow$
+SC (GPT-3.5) 37.91 (42.68$\downarrow$ 47.47 (47.21$\uparrow$ 40.12 (39.10$\uparrow$ 44.86 (47.61$\downarrow$ 47.00 (48.48$\downarrow$ 40.74 (40.63$\uparrow$
+SC (GPT-4o) 47.36 (44.78$\uparrow$ 41.17 (40.77$\uparrow$ 43.79 (45.93$\downarrow$ 49.14 (46.83$\uparrow$ 47.84 (49.38$\downarrow$ 45.68 (46.22$\downarrow$
 Non-explicit
-BTSD (GPT-3.5) 30.73 (33.10) 41.54 (38.10) 40.76 (38.78) 32.14 (31.32) 40.38 (38.55) 40.54 (41.89)
-BTSD (GPT-4o) 32.72 (35.14) 41.71 (38.55) 41.34 (42.78) 33.44 (36.12) 41.92 (39.84) 43.90 (45.70)
-SC (GPT-3.5) 31.88 (32.66) 43.41 (45.54) 37.76 (36.69) 35.46 (33.94) 44.46 (45.80) 38.10 (38.97)
-SC (GPT-4o) 36.00 (36.39) 38.27 (39.92) 38.34 (38.50) 38.11 (37.5) 42.81 (43.84) 37.54 (39.26)
+BTSD (GPT-3.5) 30.73 (33.10$\downarrow$ 41.54 (38.10$\uparrow$ 40.76 (38.78$\uparrow$ 32.14 (31.32$\uparrow$ 40.38 (38.55$\uparrow$ 40.54 (41.89$\downarrow$
+BTSD (GPT-4o) 32.72 (35.14$\downarrow$ 41.71 (38.55$\uparrow$ 41.34 (42.78$\downarrow$ 33.44 (36.12$\downarrow$ 41.92 (39.84$\uparrow$ 43.90 (45.70$\downarrow$
+SC (GPT-3.5) 31.88 (32.66$\downarrow$ 43.41 (45.54$\downarrow$ 37.76 (36.69$\uparrow$ 35.46 (33.94$\uparrow$ 44.46 (45.80$\downarrow$ 38.10 (38.97$\downarrow$
+SC (GPT-4o) 36.00 (36.39$\downarrow$ 38.27 (39.92$\downarrow$ 38.34 (38.50$\downarrow$ 38.11 (37.5$\uparrow$ 42.81 (43.84$\downarrow$ 37.54 (39.26$\downarrow$
 Table 7: Preliminary results using the CoT prompting strategy with GPT-3.5 and GPT-4o models on three datasets (TSE, VAST, EZSTANCE), evaluated for both Target Generation (TG) using BTSD and Stance Detection (SD) using SC.
-Results in the parenthesis are from Table 2 following the "Task Definition" strategy.
+Results in parentheses refer to scores reported in Table 2 using the "Task Definition" prompting strategy.
+An upward arrow ($\uparrow$) indicates that the CoT strategy outperformed the Task Definition approach, while a downward arrow ($\downarrow$) indicates a performance drop.
 The results are presented for both Explicit (Ex) and Non-explicit (N-Ex) cases across two approaches: TG+SD and TG&SD.
-Appendix J Chain-of-Thought (CoT) Results
+Appendix I Chain-of-Thought (CoT) Results
 We conduct a preliminary experiment using the CoT prompting strategy across both approaches: TG+SD and TG&SD in Table 7.
@@
 The final output format MUST be: “‘Target: <target>, Stance: <stance>’ ”.
+Appendix J Preliminary Study on Coherence Measure in OTSD
+To address the need for a more accurate evaluation metric in the Open Target Stance Detection (OTSD) task—one that aligns the generated stance with the corresponding generated target—we conduct a small-scale experiment that might be leading to a promising direction for developing a robust coherence-based measure.
+Specifically, we employ an LLM to assess how coherent a generated stance is with respect to the generated target, given the context of a text.
+For our experiment, we selected 200 samples from the EZSTANCE dataset (Zhao and Caragea, 2024), covering all configurations (i.e., Explicit and Non-explicit, TG+SD and TG&SD).
+We use LLaMA-3-8B as the evaluation model, and choose the same "Task Definition" prompting strategy as applied in Section 3.2 but with few-shot samples.
+Each prompt includes the input text, the generated target, and the corresponding generated stance.
+As this is a preliminary experiment, future work could explore a variety of alternative prompting strategies to further improve performance.
+The prompt used in our work is as follows:
+"You will be given a text, a generated stance (generated_stance), and a generated target (generated_target).
+Your task is to determine whether the stance is coherent with respect to the target in the context of the text.
+A stance is coherent if it correctly reflects the text’s position towards the target as implied or stated in the text.
+Analyze the text well, as it may convey the stance towards the target implicitly.
+Output 1 if the stance is coherent, and 0 if it is not.
+Output only a single digit: 0 or 1, with no explanation.
+Input Format:
+text: <tweet text>
+generated_target: <target>
+generated_stance: <stance>
+Examples:
+text: "Renewable energy is the only viable path forward if we care about this planet."
+generated_target: renewable energy
+generated_stance: AGAINST
+Output: 0
+text: "Elon Musk is brilliant, but his take on public transport is just awful."
+generated_target: Elon Musk
+generated_stance: FAVOR
+Output: 1
+text: "Vaccines have done more good than harm. I trust the science."
+generated_target: vaccines
+generated_stance: NONE
+Output: 0
+text: "As you have been told already what Phil says is irrelevant to ID.
+Or you have to allow what Dawkins says about the theory of evolution.What part of that don’t you understand?And it is a fact that if you could support your position with real data ID would go away.
+You can whine all you want but that will not change that fact."
+generated_target: Intelligent Design
+generated_stance: AGAINST
+Output: 1
+text: "I’ve seen hundreds of tweets at this point, many from people with big platforms, condemning leftists who are apparently pro-Putin.
+I have yet to see any leftists actually express support for Putin.
+All I’ve seen from the left is unambiguous solidarity for the people of Ukraine."
+generated_target: leftists
+generated_stance: NONE
+Output: 1"
+To further validate the LLM-based evaluation, we also collected human judgments on the same 200 samples.
+A single human annotator was asked to assess the coherence of the generated stances with respect to the generated targets, using the same task definition provided to the LLM.
+The annotator received the same input triplet: text, generated target, and generated stance.
+We then measured the correlation between the LLM’s coherence judgments and those of the human annotator using the Phi coefficient.
+The resulting score was 0.6541, indicating a strong positive correlation (Njuka and Phiri, 2021).
+It is important to note that we do not propose a definitive solution or formal methodology for coherence evaluation in stance detection.
+Rather, we present this as a potential path that can be further explored by the research community.
+Our experiment suggests that LLMs might serve as reliable judges for assessing the coherence between a generated stance and target within a given text, with judgments that align well with human evaluations.
+To strengthen the validity of this experiment, future studies could incorporate a wider range of LLMs, multiple runs per sample, and a larger, more diverse set of samples across the three datasets.
 Appendix K Human Annotation Score Distribution
@@
 In cases of a tie, the average score of 0.5 is assigned.
-The details on human annotator guideline are provided in Appendix I.1, and the annotator agreement is in Appendix Table 6.
+The details on human annotator guideline are provided in Appendix H.1, and the annotator agreement is in Appendix Table 6.